### v1

# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

Abstract
Conventional survival metrics, such as Harrell’s concordance index and the Brier Score, rely on the independent censoring assumption for valid inference in the presence of right-censored data. However, when instances are censored for reasons related to the event of interest, this assumption no longer holds, as this kind of dependent censoring biases the marginal survival estimates of popular nonparametric estimators. In this paper, we propose three copula-based metrics to evaluate survival models in the presence of dependent censoring, and design a framework to create realistic, semi-synthetic datasets with dependent censoring to facilitate the evaluation of the metrics. Our empirical analyses in synthetic and semi-synthetic datasets show that our metrics can give error estimates that are closer to the true error, mainly in terms of predictive accuracy.
1 Introduction
Learning survival prediction models resembles learning regression models, as both are essentially trying to learn, from a labeled dataset, a model that maps a description of an instance to a real value. However, survival prediction differs as some of the training instances are “censored” – that is, we only know a lower bound of the true time. This includes, for example, individuals who are still alive at the end of the study, as well as others who left the study and were “lost to follow-up”. The independent censoring assumption states that, within any subgroup of interest, we assume that the patients (i.e., instances) who are censored at time $t$ are representative of all the patients in that subgroup who remained at risk at time $t$ with respect to their survival experience [Emura and Chen, 2018, Ch. 1]. That is, censoring is random within the subgroups. Let $T_{E}$ be the event distribution and $T_{C}$ be the censoring distribution, then we have $T_{E}\perp T_{C}$i.e.$T_{E}$ provides no information about the distribution of $T_{C}$, and vice versa [Kleinbaum and Klein, 2012, Ch. 1]. If there are features (covariates) $X$, this assumption becomes the conditional independent censoring assumption, i.e. $T_{E}\perp T_{C}\mid X$ [Kleinbaum and Klein, 2012, Ch. 1]. (More details wrt censoring are in Appendix C.) Note that many standard survival tools – e.g., the Kaplan-Meier (KM) [Kaplan and Meier, 1958] and Nelson-Aalen [Nelson, 1969, Aalen, 1978] estimators, and the semiparametric Cox Proportional Hazards (CoxPH) model [Cox, 1972], rely on an assumption of independent censoring for valid inference in the presence of right-censored data [Kleinbaum and Klein, 2012, Ch. 1].
Figure 1: (a) Residual dependence between event time $T_{E}$ (cancer relapse) and censoring time $T_{C}$ (cancer death) denoted by the diamond ‘Dep’, due to an unobserved confounding covariate – here ‘Tumor grade’. (b) Evaluation under dependent censoring, where Patient A is censored immediately before the event for reasons related to the event. This leads to a bias estimation on the marginal survival distribution using the Kaplan-Meier (KM) [Kaplan and Meier, 1958] estimator. Note that the two curves are coincide between range $[0,t_{c,i}]$. (c) Scatter plot generated from a standard exponential distribution under a Clayton [Clayton, 1978] copula model.
Figure 1(a) shows how the independent censoring assumption fails to hold by omitting a confounding covariate – e.g.“tumor grade”. Since “tumor grade” relates to both $T_{E}$ (cancer relapse) and $T_{C}$ (cancer death), the variation in the covariate induces changes in both $T_{E}$ and $T_{C}$. That is, we want a model that predicts the time until relapse, but death is censoring. As “tumor grade” is predictive of tumor progression, an advanced (resp., early) stage is linked to shorter (longer) values for both cancer relapse $T_{E}$ and cancer death $T_{C}$, and consequently, $T_{E}$ and $T_{C}$ are not independent. This shows that the conditional independent censoring assumption might not hold if any confounding covariates are omitted or ignored in a regression model [Emura and Chen, 2018, Ch. 1]. This phenomenon is called residual dependencei.e.dependent censoring not adjusted by the covariates.
Figure 1(b) demonstrates how dependent censoring biases the marginal survival estimates from the KM estimator [Kaplan and Meier, 1958]. Consider patient A, censored (due to cancer death) exactly one year after study entry, who likely would have experienced cancer relapse shortly thereafter. However, the KM estimator may overestimate this patient’s survival time by assuming that they are like those remaining in the study, who tend to live longer, because it does not consider that these censored patients tend to have shorter event times. Indeed, the estimator relies on a stepwise construction of the survival probability that depends on knowing how many individuals are at risk and the event probabilities at each time point [Bland and Altman, 1998].
To address such dependency, Figure 1(c) presents a copula-based modeling approach. A copula function $C_{\theta}(u,v)$ links random variables by specifying their dependence structure [Emura and Chen, 2018, Ch. 1]. Applied here, it captures the dependency between event and censoring distributions ($T_{E}$ and $T_{C}$) by sampling the marginals from a (survival) copula. In this example, we see lower tail dependence from the copula – describing the probability that small (early) values of one variable (e.g., event) are associated with small values of another variable (e.g., censoring).
In the absence of features, if any confounder related to the event censors the event time, we refer to this phenomenon as dependent censoring. However, it is typically impossible to verify this in practice because we only observe one outcome (either event or censoring) per instance, but not both [Tsiatis, 1975]. Although we can reduce dependent censoring simply by collecting more data (i.e., features) [Emura and Chen, 2018, Ch. 1], we would still have to evaluate our model under dependent censoring, independent of the learned model, preferably. Popular metrics for evaluating predicted survival probabilities (or risk scores) are Harrell’s concordance index (CI) [Harrell Jr et al., 1982], Uno’s CI [Uno et al., 2011], the Brier Score (BS) [Brier, 1950], and the mean absolute error (MAE) using one of hinge [Haider et al., 2020], margin [Haider et al., 2020] or pseudo times [Qi et al., 2023a]. Unfortunately, dependent censoring is both common, and also introduces a bias that can affect the estimation of survival probabilities, and none of these metrics account for this dependency. Alternative methods have been proposed to handle dependent censoring, such as the inverse probability of censoring weighting (IPCW) [Robins and Finkelstein, 2000], but this approach still uses the KM estimator. Motivated by these limitations, we introduce a new approach to evaluating survival models under dependent censoring. To our knowledge, no prior work has proposed evaluation metrics that explicitly model the degree of dependent censoring between $T_{E}$ and $T_{C}$. Our main contributions are:
  • We show that conventional survival metrics, such as Harrell’s CI, Uno’s CI, the IPCW-integrated Brier Score, and the marginal MAE, are biased under dependent censoring, meaning they will give inaccurate estimates of the model’s discriminative or predictive performance.
  • We propose three novel metrics – CI-Dep, IBS-Dep, and MAE-Dep – to evaluate survival models in the presence of dependent censoring. These metrics can evaluate the concordance (i.e., ranking), the accuracy of the survival curve, and the accuracy of the predicted time-to-event estimates, respectively, by considering the dependency between the censoring and event distributions.
  • We design a framework to create realistic, semi-synthetic datasets with various degrees of dependent censoring to determine if the proposed metrics can (1) report an unbiased error closest to the true error, and (2) rank the survival models by the true error.
  • We provide empirical evidence that our proposed metrics can provide more accurate error measures for survival models, mainly in terms of predictive accuracy, i.e., Brier Score and MAE.
Our code is publicly available at this GitHub repository.
2 Background and related work
2.1 Survival data and notation
Many survival analysis models estimate the probability that a specific event will occur for an individual at a time $T$ later than $t$, for all times $t>0$. Examples include predicting the time to death in patients with amyotrophic lateral sclerosis (ALS) [Kuan et al., 2023], predicting the time to a fall in older adults [Lillelund et al., 2024a], or evaluating a new experimental treatment for cancer [Zhang et al., 2011]. Most often, the observation for a given instance will consist of a random variable, $T$, representing the time from a given origin ($t=0$) to the occurrence of the event, e.g.“death”. The distribution of $T$ can be characterized by the probability distribution function $F(t)=\text{Pr}(T\leq t)$ or, equivalently, by the survival function $S(t)=1-F(t)=\text{Pr}(T>t)$. Note that $S(t)$ (resp., $F(t)$) correspond to the probabilities of being event-free (resp., having experienced the event by) time $t$. We can estimate $S\left(t\right)$ nonparametrically, e.g., using the Kaplan-Meier [Kaplan and Meier, 1958] estimator (see Appendix H for a detailed derivation).
Definition 1
The Kaplan-Meier (KM) estimator
$\begin{split}\hat{S}(t)\ &=\prod_{t_{i}\leq t,d_{i}=1}\left(1-\frac{1}{n_{i}}\right),\quad\hbox{\rm for}\quad t\leq t_{\text{max}}\end{split}$ (2.1)
where $n_{i}=\sum_{\ell=1}^{n}I(t_{\ell}\geq t_{i})$ is the number at-risk at time $t_{i}$; $\hat{S}(t)=1$ if no death occurs up to time $t$; $\hat{S}(t)$ is undefined for $t>t_{\text{max}}$ [Kaplan and Meier, 1958].
If features are entered into the model, popular methods include the CoxPH model [Cox, 1972, Katzman et al., 2018, Lillelund et al., 2024b], gradient-boosted regression [Ridgeway, 1999], Random Survival Forest [Ishwaran et al., 2008] and Multi-Task Logistic Regression (MTLR) [Yu et al., 2011]. In this context, we may define our survival dataset $\mathcal{D}$ as a set of observations, i.e., we have $\mathcal{D}=\left\{\left(\bm{x}_{i},t_{i},\delta_{i}\right)\right\}_{i=1}^{N}$, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a feature (covariate) vector for instance $i$, $t_{i}\in\mathbb{R}^{\geq 0}$ is the time of either censoring or event, depending on which occurred first, and $\delta_{i}\in\{0,1\}$ is the “censoring bit”: If $\delta_{i}=0$, then $t_{i}=t_{c,i}$, where $t_{c,i}$ is the time of censoring, otherwise, if $\delta_{i}=1$, then $t_{i}=t_{e,i}$, where $t_{e,i}$ is the time of the event. Given survival data $\mathcal{D}$, let $f_{T\mid X}(\cdot)$ and $F_{T\mid X}(\cdot)$ represent the conditional density and cumulative distribution functions, respectively, over the event horizon (i.e., from 00 to $t_{\text{max}}$). Given the covariates, we then have the following useful definitions:
Definition 2
The survival function
$S_{T\mid X}(t\mid X)=\text{Pr}(T>t\mid X)=1-F_{T\mid X}(t\mid X),$ (2.2)
represents the probability the event occurs at time $t$ [Gareth et al., 2021, Ch. 11].
Definition 3
The hazard function
$h_{T\mid X}(t\mid X)$ =\ = $\lim_{\epsilon\to 0}\text{Pr}(T\in[t,t+\epsilon)\mid T\geq t,X)$ (2.3)
=\ = $\frac{f_{T\mid X}(t\mid X)}{S_{T\mid X}(t\mid X)}$
represents the instantaneous event rate at given time $t$, conditional on surviving $t$ [Gareth et al., 2021, Ch. 11].
2.2 Copula-Based Estimation
A copula is a function that links two or more random variables by specifying their dependence structure [Emura and Chen, 2018, Ch. 1]. Among the various types of copulas, the Archimedean copula family is the most used, and enables modeling of dependencies in arbitrarily high dimensions using a single parameter, $\theta$. This family of copulas is defined by a generator function, $\varphi_{\theta}(t):[0,1]\to[0,\infty]$, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [Emura and Chen, 2018, Ch. 3]. Here, we will focus on two common Archimedean copulas: Clayton [Clayton, 1978] and Frank [Frank, 1979].
Zheng and Klein [1995] generalized the KM estimator under independent censoring to the Copula-Graphic (CG) estimator under dependent censoring, which has become an important tool for analyzing survival data [Emura and Chen, 2016, Emura and Michimae, 2017, Hoora Moradian and Bellavance, 2019]. Rivest and Wells [2001] obtained the first explicit expression of the CG estimator under an assumed copula from the Archimedean family (e.g., Clayton or Frank). Given survival data without covariates, i.e.$(t_{i},\delta_{i})$ for $i=1,\ldots,n$, where $t_{i}=\min(t_{e,i},t_{c,i}),\;\delta_{i}=I(t_{e,i}\leq t_{c,i})$ with $I(\cdot)$ denotes the indicator function, we can estimate the survival function using the CG estimator (see Appendix H for a detailed derivation):
Definition 4
The Copula-Graphic (CG) estimator
$\small\begin{split}\hat{S}_{T}(t)&=\varphi^{-1}\left(\sum_{{i:\,}t_{i}\leq t,\delta_{i}=1}\varphi\left(\frac{n_{i}-1}{n}\right)-\varphi\left(\frac{n_{i}}{n}\right)\right),\\ \hbox{for}&\quad t\ \leq\ t_{\text{max}}\end{split}$ (2.4)
where $n_{i}=\sum_{\ell=1}^{n}I(t_{\ell}\geq t_{i})$ is the number at risk at time $t_{i}$; $\hat{S}_{T}(t)=1$ if no death occurs up to time $t$; $\hat{S}_{T}(t)$ is undefined for $t>t_{\text{max}}$ [Emura and Chen, 2018, Ch. 4].
Note that when $\varphi(\cdot)$ is uniform, this reduces to the standard KM estimator. See Appendix D for more details.
2.3 Evaluation metrics and limitations
We now outline the conventional evaluation metrics in survival analysis. The detailed calculation of these metrics is outlined in Appendix E and here we mainly focus on why those metrics are problematic under dependent censoring.
Concordance index (CI) evaluates the ranking of individuals when their relative ordering is of primary interest. It measures the proportion of concordant pairs – pairs where the predicted ranking of risk scores aligns with the true ordering of event times – among all comparable pairs [Harrell Jr et al., 1996]. A pair $(i,j)$ is defined as comparable if $\delta_{i}=1$ and $t_{i}<t_{j}$, which means that the earlier individual has experienced the event and the observed time of the later individual is greater than the first. Even under the random censoring assumption, it is evident that the comparable pairs selected using the above criteria do not represent all possible pairs based on their true event times:
          $\text{Pr}(\delta_{i}=1,t_{i}<t_{j})\quad\neq\quad\text{Pr}(t_{e,i}<t_{e,j}).$
This inequality arises because the left-hand term can be decomposed as:
$\begin{array}[]{l}\text{Pr}(\,\delta_{i}=1,\ t_{i}<t_{j})\\ =\quad\text{Pr}(\,t_{e,i}<t_{c,i},\ \ t_{e,i}\ <\ \min\{t_{c,j},t_{e,j}\}\ )\\ =\quad\text{Pr}(\,t_{e,i}<t_{c,i},\ t_{e,i}<t_{c,j},\ t_{e,i}<t_{e,j})\\ =\quad\underline{\text{Pr}(\,t_{e,i}<t_{c,i})\ \text{Pr}(t_{e,i}<t_{c,j})}\quad\text{Pr}(t_{e,i}<t_{e,j})\ ,\end{array}$
where the factorization follows from the independent censoring assumption. Compared to the right-hand side term, this expression includes additional underlined terms related to the censoring distribution, confirming the findings of Uno et al. [2011].
To address this limitation, Uno et al. [2011] introduced a modified version of CI that employs inverse probability of censoring weights (IPCW), ensuring consistency with the true CI. The IPCW approach estimates the probability of being uncensored at each time point and reweights the uncensored instances accordingly.
Integrated Brier Score (IBS) evaluates the overall accuracy of predicted survival functions by integrating the Brier Score (BS) [Brier, 1950] over a range of time points. It measures the average squared difference between the predicted survival probabilities and actual survival outcomes over time [Gerds and Schumacher, 2006]. To handle censored instances, IPCW can be used to reweight uncensored instances when computing both BS and IBS [Graf et al., 1999]. Similar to IPCW-CI, IPCW weights are often approximated using the KM estimator, which is biased under dependent censoring [Emura and Chen, 2018, Ch. 1].
Mean Absolute Error (MAE) quantifies the accuracy of point predictions for event times by measuring the absolute difference between the predicted and true event times. For censored instances, Haider et al. [2020] proposed an approach that approximates that subject’s event time as the mean of the KM estimator (Equation 2.1), renormalized at that individual’s censoring time. This proxy event time, weighted by confidence estimates (also derived from KM), is then used to compute the absolute error. We will refer to this method as MAE-Margin. Given a subject is censored at time $t_{i}$, its margin time is calculated as:
[…]
where $\hat{S}(t)$ is an estimator, e.g., the KM, derived from the training dataset.
Limitations. All the aforementioned metrics typically rely on KM estimates to approximate either the marginal time-to-event or the marginal time-to-censoring distribution. While KM is advantageous due to its consistency under the independent censoring assumption and its computational simplicity – requiring no interaction with covariates – it becomes problematic when censoring and event times are dependent [Emura and Chen, 2018, Ch. 1]. In such cases, the KM estimates may be biased and inconsistent [Campigotto and Weller, 2014]. To mitigate these issues, researchers have proposed modifying these metrics by replacing KM with a conditional estimator (e.g.,  CoxPH) to account for the dependency between covariates and censoring distributions [Gerds and Schumacher, 2006]. However, this approach still has two major limitations: (1) model misspecification: if the conditional estimator is inaccurately specified, the resulting estimates may be unreliable, and (2) uncaptured dependencies: under dependent censoring (i.e.,  when unobserved confounders exist), even a well-specified conditional estimator may fail to fully account for the dependency if there are important covariates missing in the analysis. These challenges highlight the need for robust estimation methods that can better handle dependent censoring scenarios.
3 The Proposed Metrics
CI-Dep: We propose a modified version of Harrell’s CI [Harrell Jr et al., 1982] that employs inverse probability weighting (i.e., IPCW), but the marginal survival probability for censoring at time $t$ is estimated using the CG estimator under an approximated Archimedean copula, parameterized by $\theta$. When $\theta\approx 0$, it yields the independence copula, which is equivalent to using the KM estimator as in Uno et al. [2011]. Our proposed concordance index thus requires a known copula, that models the dependency between marginals $T_{E}$ and $T_{C}$, but does not require interaction with the covariates. Given an estimated risk function $\hat{\eta}(\cdot)$ and corresponding labels, the traditional CI is computed as:
$\small\begin{split}\text{CI}(\mathcal{D};\,\hat{\eta}(\cdot))=\frac{\sum_{i,j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]}.\end{split}$ (3.1)
Now, employing the IPCW technique used by Uno et al. [2011], but under dependent censoring, let $\hat{G}_{\text{CG}}(t)=\text{Pr}(T>t)$ represent the marginal survival probability for censoring at $t$ estimated by the CG estimator. We can then reweight Equation 3.1 using the CG estimator as:
$\small\begin{split}&\text{CI}_{\text{Dep}}(\mathcal{D};\,\hat{\eta}(\cdot))\ =\\ &\frac{\sum_{i,j}\delta_{i}\cdot\{\hat{G}_{\text{CG}}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j}\delta_{i}\cdot\{\hat{G}_{\text{CG}}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]},\end{split}$ (3.2)
which denotes the proposed CI-Dep. Here, the weight is given by $1/\hat{G}_{\text{CG}}(t)$.
IBS-Dep: We propose a modified version of the Brier Score [Graf et al., 1999] under dependent censoring using margin time estimates as the censoring weights. The margin or “best-guess” time can be interpreted as a conditional expectation of the event time given the event time is greater than the censoring time. Given an instance is censored at time $t_{i}$, we can calculate its margin time using Equation 2.5, but with the CG estimator instead of the KM, i.e.$t_{e}^{\text{margin}}(t_{i})=t_{e}^{\text{margin}}(\hat{S}_{CG},t_{i})$, where $\hat{S}_{CG}$ is the marginal survival probability estimated using CG. We then formulate the dependent Brier Score as:
$\begin{array}[]{l}\text{BS}_{\text{Dep}}\left(\mathcal{D};\,\hat{S}(\cdot),t^{*}\right)\\ =\frac{1}{N}\sum_{i=1}^{N}\Bigg{(}\mathbbm{1}[\tilde{e}_{i}>t^{*}]-\hat{S}(t^{*}\mid\bm{x}_{i})\Bigg{)}^{2}.\end{array}$ (3.3)
$\text{where}\quad\tilde{e}_{i}=\begin{cases}t_{i}&\text{if}\quad\delta_{i}=1\\ t_{e}^{\text{margin}}(t_{i})&\text{elsewhere.}\end{cases}$
To calculate the expectation of Equation 3.3 over all times, we simply take the integral from $t=0$ to $t=t_{\text{max}}$
[…]
where $t_{\text{max}}$ is usually the maximum observed time.
MAE-Dep: We propose a modified version of the MAE-Margin [Haider et al., 2020] under dependent censoring. Again, we make use of the CG estimator to calculate margin times for the censored instances (Equation 2.5) and replace the censored event times with their “best-guess”-estimates. Furthermore, for each censored instance, we adopt the approach from Haider et al. [2020] and assign a confidence weight $\omega_{i}=1-S_{\text{\text{CG}}}(t_{c,i})$ to the error based on the margin value estimated using the CG method. For uncensored subjects, we set $w_{i}=1$. The dependent MAE is thus:
$\text{MAE}_{\text{Dep}}\left(\mathcal{D};\,\hat{t}_{i}\right)=\frac{1}{\sum_{i=1}^{N}\omega_{i}}\sum_{i=1}^{N}\omega_{i}\Big{|}\big{[}(1-\delta_{i})\cdot t_{e}^{\text{margin (CG)}}(t_{i})$
$\qquad+\quad\delta_{i}\cdot t_{i}\big{]}-\hat{t}_{i}\Big{|}\ .$ (3.5)
4 Experiments and results
For our empirical analyses, we evaluate the effectiveness of the proposed metrics under dependent censoring in synthetic datasets under a known copula and in semi-synthetic datasets by learning the copula based on the data.
4.1 Experimental Setup
Figure 2: Outline of the proposed evaluation pipeline. Based on some raw, survival dataset $\mathcal{D}$, we (1) form $\tilde{\mathcal{D}}$ by flipping the event bit of $\mathcal{D}$ and then estimate the feature-dependent censoring distribution, $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$. We then (2) select a subset of features from $\mathcal{D}$ using some strategy (e.g., Top-K features), and (3) generate a ground-truth dataset, $\mathcal{D}^{\prime}$, by removing the censored instances from $\mathcal{D}$. We then (4) create $\mathcal{D}^{\prime\prime}$ based on the ground-truth dataset, the learned censoring distribution, and the selected features, $\{X_{j}\}_{j\in\mathcal{S}}$. Using $\mathcal{D}^{\prime\prime}$, we (5) train two copula models, i.e., Clayton and Frank, and use AIC to select the copula $C_{\theta}$ that best suits the data. At the same time, we (6) train 5 survival learners independent of the copula on $\mathcal{D}^{\prime\prime}$ to produce survival predictions. Lastly, we calculate the (censored) evaluation error and (7) compare this to the true error.
Synthetic datasets. Following the approach of Foomani et al. [2023], we create a synthetic survival dataset from a data generation process (DGP). The event and censoring distributions are specified by linear Weibull marginal distributions. Specifically, we sample $T_{E}$ and $T_{C}$ random variables from a known copula given Kendall’s $\tau$; if $\tau=0$, the marginals are independent, otherwise if $\tau>0$, the marginals are from an Archimedean copula, e.g., Clayton or Frank, known at evaluation time. See Appendix F.1 for more details.
Semi-synthetic datasets. To evaluate our proposed metrics, we need to know the true metric, which requires precise knowledge of when each instance experiences the event. Due to censoring, event times are not available in real-world datasets for all instances, so we create a realistic, semi-synthetic dataset, where the covariates, event distribution, and censoring distribution align with some real-world datasets. This approach is inspired by Qi et al. [2023a], who proposed a semi-synthetic framework to generate different kinds of censoring distributions. However, in our case, we propose several strategies instead to create dependent censoring. In summary, our approach is as follows (see also the flowchart in Figure 2):
  1. 1.
    Start with some real-world survival dataset $\mathcal{D}$.
  2. 2.
    Based on some feature selection strategy (e.g., Top-$k$ features), select the relevant features from the original dataset $\mathcal{D}$ based on some technique, e.g., feature importances.
  3. 3.
    Calculate the censoring distribution of $\mathcal{D}$ to generate censor times, $\tilde{t}_{c,i}\sim G_{\text{CoxPH}(\tilde{\mathcal{D}})}(t\mid x_{i}$), where $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$ is the feature-dependent censoring distribution estimated by the Cox Proportional Hazard model [Cox, 1972] with Breslow estimator [Breslow, 1975].
  4. 4.
    Create $\mathcal{D}^{\prime}$ by removing the censored instances from $\mathcal{D}$, leaving just the uncensored ones.
  5. 5.
    Form $\mathcal{D}^{\prime\prime}$ by applying some reasonable, but synthetic censoring distribution to $\mathcal{D}^{\prime}$ and restrict the feature set to induce dependent censoring. Here we use the censoring distribution calculated in step 3 and the features calculated in step 2.
Based on $\mathcal{D}$, we calculate the censoring distributions $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$ and apply it to $\mathcal{D^{\prime}}$, so that for each instance, that instance is synthetically censored if the censoring time is earlier than the event time ($\tilde{t}_{c,i}<t_{i}$), and otherwise is left uncensored. To induce dependent censoring, we consider three different feature selection strategies:
  • Original, where the original feature set is kept intact. This approach does not introduce additional residual dependence, but conditional independence may still not hold if important features are omitted originally.
  • Top-$k$, where we select the $k$ most important features from $\mathcal{D}$ based on their predictive power. For this, we train a CoxPH model on $\mathcal{D}$ and then use permutation feature importances [Breiman, 2001] to select the $k$ best features. Feature importances are calculated by shuffling feature values and observing how much the model’s performance degrades, using Harrell’s CI.
  • Random-%, where we select a random proportion of %-pct features from the original feature set. In a small dataset, randomly selecting 25% of the features can introduce significant residual dependence if key features are excluded. In contrast, in a high-dimensional dataset, this risk is lower as many important features are likely to remain.
Copula model. To validate our dependent metrics in semi-synthetic datasets, we train two copula models, Clayton and Frank, and two nonlinear Weibull proportional-hazards models for the marginal distributions, $T_{E}$ and $T_{C}$, respectively, on $\mathcal{D}^{\prime\prime}$. The Weibull models use a multilayer perceptron (MLP) as the backbone network architecture with a single hidden layer and the nonlinear ReLU activation function. For each copula model, we train the marginals and the copula $C_{\theta}$ jointly using maximum likelihood estimation (MLE) (see Appendix F.3 for details). After training the two candidate models, we choose the model that minimizes the Akaike Information Criterion (AIC), given its loss and number of parameters. This offers a practical, data-driven approach to copula selection [Emura and Chen, 2018, Ch. 4]. We then use the best copula $C_{\theta}$ for the evaluation of the dependent metrics.
Datasets. We apply the evaluation pipeline in Figure 2 to 6 large, real-world datasets: METABRIC [Curtis et al., 2012], 2 from MIMIC-IV (all causes and hospital death) [Johnson et al., 2023], and 3 from SEER (death from brain, liver and stomach cancer) [LAG et al., 2007]. They differ in their number of instances, features, and censoring rates. Appendix F.1 gives further details on the datasets and describes how we preprocessed them.
Survival learners. We compare the prediction results using 5 survival learners produced by the CoxPH [Cox, 1972], gradient boosting (GBSA) [Ridgeway, 1999], Random Survival Forest (RSF) Ishwaran et al. [2008], DeepSurv Katzman et al. [2018] and Multi-Task Logistic Regression (MTLR) [Yu et al., 2011]. These learners predict individual survival distributions (ISDs) for each instance $i$. See Appendix F.2 for further details, including implementation details and hyperparameters.
Figure 3: Plot of the bias – i.e., the proximity to the true error – of established baseline metrics and the proposed dependent metrics as a function of the dependence under a known copula. We plot the mean biases across 10 experiments using different random seeds with data generated from a DGP ($N=10,000$, $d=10$). Overall, our dependent metrics show lower biases when dependency is present ($\tau>0.25$) than the independent metrics. See Appendix F.1 for details.
Given the semi-synthetic dataset $\mathcal{D}^{\prime\prime}$, we train the five survival learners across 5 experiments given a random seed. In each experiment, we follow Sechidis et al. [2011] and first split the dataset into training (70%), validation (10%), and test (20%) sets in a stratified manner wrt the event indicators and event times. This ensures a balanced representation of the time-to-event information and outcome across the data splits. After splitting, we impute missing values using sample mean for real-valued covariates or mode for categorical covariates based on the training set only. We then normalize numerical values and encode categorical features using a one-hot encoding strategy.
4.2 Results and Takeaways
Synthetic data. Figure 3 shows the evaluation bias – i.e., the proximity to true error – for four baseline metrics (Harrell’s CI, Uno’s CI, IPCW-integrated Brier Score, and MAE-Margin) and their dependent variants as a function of Kendall’s $\tau$. Assuming a known ground-truth copula (Archimedean family and dependency strength), we estimate $\hat{S}(t)$ using a CoxPH model to predict $N$ ISDs, where the risk scores are the negative predicted survival times. Overall, our dependent metrics show lower bias – i.e., closer proximity to the true error – when dependency is present ($\tau>0.25$). For the ranking metrics (Harrell’s CI, Uno’s CI), biases are similar under low dependency ($\tau<0.25$) but improve significantly under high dependency ($\tau>0.5$). For accuracy metrics (IPCW-integrated Brier Score, MAE-Margin), our dependent variants consistently provide better error estimates regardless of dependency or copula family. Discrepancies are most notable under a Clayton copula with moderate dependency ($0.50<\tau<$ 0.5).
As a practical example, consider a cancer study, with high dependent censoring ($\tau=0.75$) from a Clayton copula between event time $T_{E}$ (cancer relapse) and censoring time $T_{C}$ (cancer death). Based on Figure 3, if the true CI of a CoxPH model is 0.65, then, if we evaluate the model using Harrell’s CI, we anticipate the CI to be 0.685 ($\pm$ 0.035). If we use CI-Dep, we anticipate the CI to be 0.675 ($\pm$ 0.025).
Semi-synthetic data. Table 1 presents the evaluation biases in 3 selected datasets: METABRIC, MIMIC-IV and SEER. The complete results for all 6 datasets are in G.1. In the small-sized METABRIC dataset, Harrell’s gives superior estimates of the true concordance index, and there is a performance degradation in using the dependent variant. We see that CI-Dep degrades the performance compared to Uno’s CI, in this case, while IBS-Dep improves on the IPCW version, and MAE-Margin and MAE-Dep give similar results. In the medium-sized MIMIC-IV dataset, Harrell’s CI is again better, but our IBS-Dep and MAE-Dep improve on the IBS-IPCW and margin MAE, respectively, in all cases, however, the pseudo MAE is the best (i.e., it has the lowest error). In the large-sized SEER dataset, there are few differences in discriminative performances for the three metrics, but our IBS-Dep improves on IPCW, while our MAE-Dep improves on pseudo and margin. More often than not, Harrell’s and Uno’s CI (IPCW) give better estimates than our CI-Dep in these datasets, while our IBS-Dep and MAE-Dep give notable improvements to their independent counterparts. See Appendix G.1 for additional results.
Dataset Strategy Discrimination (CI) Accuracy (IBS) Accuracy (MAE)
Harrell’s Uno’s Dependent IPCW Dependent Uncens. Hinge Pseudo Margin Dependent
METABRIC ($N$=1,102, $C$=3.8%) Original 0.10$\pm$0.08 0.39$\pm$0.14 0.48$\pm$0.25 1.23$\pm$0.85 1.17$\pm$0.85 3.24$\pm$0.55 1.40$\pm$0.21 0.25$\pm$0.21 0.31$\pm$0.23 0.33$\pm$0.22
Top 5 0.11$\pm$0.08 0.51$\pm$0.17 0.58$\pm$0.23 1.22$\pm$0.85 1.17$\pm$0.85 3.21$\pm$0.55 1.40$\pm$0.21 0.25$\pm$0.20 0.31$\pm$0.23 0.30$\pm$0.23
Top 10 0.10$\pm$0.08 0.43$\pm$0.11 0.49$\pm$0.16 1.23$\pm$0.85 1.17$\pm$0.85 3.26$\pm$0.54 1.40$\pm$0.21 0.25$\pm$0.21 0.32$\pm$0.23 0.31$\pm$0.23
Rand. 25% 0.10$\pm$0.08 2.34$\pm$0.68 2.31$\pm$0.70 1.25$\pm$0.88 1.18$\pm$0.88 3.45$\pm$0.48 1.40$\pm$0.21 0.27$\pm$0.22 0.34$\pm$0.23 0.33$\pm$0.23
MIMIC-IV (hospital) ($N$=6,780, $C$=74.3%) Original 0.74$\pm$0.43 1.27$\pm$0.83 2.10$\pm$1.54 8.75$\pm$1.50 6.57$\pm$1.44 3.48$\pm$0.25 6.44$\pm$0.27 3.11$\pm$0.27 4.54$\pm$0.26 4.36$\pm$0.24
Top 5 0.79$\pm$0.55 1.16$\pm$0.56 3.36$\pm$2.74 8.80$\pm$1.51 6.42$\pm$1.45 3.51$\pm$0.24 6.45$\pm$0.26 3.10$\pm$0.26 4.65$\pm$0.25 4.33$\pm$0.18
Top 10 0.71$\pm$0.47 1.10$\pm$0.65 2.46$\pm$1.69 8.73$\pm$1.50 6.57$\pm$1.44 3.50$\pm$0.27 6.45$\pm$0.27 3.11$\pm$0.27 4.60$\pm$0.28 4.40$\pm$0.25
Rand. 25% 1.44$\pm$0.58 1.77$\pm$1.02 2.70$\pm$1.44 10.38$\pm$0.76 7.91$\pm$0.90 3.57$\pm$0.20 6.51$\pm$0.19 3.06$\pm$0.19 4.79$\pm$0.19 4.62$\pm$0.20
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.77$\pm$0.10 0.57$\pm$0.14 0.76$\pm$0.13 0.33$\pm$0.15 0.25$\pm$0.11 2.59$\pm$0.07 2.50$\pm$0.04 2.47$\pm$0.07 1.48$\pm$0.06 1.44$\pm$0.05
Top 5 1.86$\pm$0.12 1.70$\pm$0.15 1.76$\pm$0.13 0.28$\pm$0.14 0.47$\pm$0.21 2.80$\pm$0.05 2.69$\pm$0.06 2.25$\pm$0.07 1.25$\pm$0.06 0.85$\pm$0.47
Top 10 0.81$\pm$0.08 0.90$\pm$0.09 1.10$\pm$0.07 0.33$\pm$0.15 0.25$\pm$0.11 2.61$\pm$0.08 2.51$\pm$0.04 2.46$\pm$0.06 1.47$\pm$0.05 1.43$\pm$0.05
Rand. 25% 0.89$\pm$0.52 1.80$\pm$1.23 1.68$\pm$1.21 0.45$\pm$0.21 0.35$\pm$0.17 2.75$\pm$0.11 2.49$\pm$0.08 2.50$\pm$0.14 1.49$\pm$0.09 1.47$\pm$0.11
Table 1: Comparison of established survival metrics and the proposed dependent metrics in terms of the mean ($\pm$ SD.) bias, i.e., the proximity to the true error, averaged over 5 learners and 5 experiments. CI and IBS results are multiplied by 100.
In general, a (censored) metric should be representative of the true metric, i.e., it should tell us which model makes the best survival predictions – for example, by identifying the top-3 learners in alignment with the true ranking. Figure 4 shows the number of times the resp. metric correctly (green dot) or incorrectly (red dot) identified the ordering of the top-3 survival learners by the true metric, over 5 experiments in the METABRIC dataset. We see general alignment between the metrics, but CI-Dep and IBS-Dep again show improvements over their independent variants.
Figure 4: Plot the ranking accuracy, i.e., the number of times the resp. metric correctly (green dot) or incorrectly (red dot) identified the true ordering of the top-3 survival learners, out of 5 experiments in the METABRIC dataset.
5 Discussion
How Well Do The Metrics Work? As we do not have a formal proof of identifiability – i.e., we do not know the true relationship between $T_{E}$ and $T_{C}$ – we therefore study this question empirically by approximating $C_{\theta}$ and making a comparison between conventional survival metrics and our dependent variants. While the advantages to adopting our dependent metrics are evident in synthetic data, where we know the ground-truth copula and dependency level, these advantages vary with the dataset and its underlying distribution in the semi-synthetic datasets. Based on our observations, there is little to no improvement in adjusting the IPCW estimates using the CG estimator for discrimination (i.e., ranking) tasks. Concerning prediction tasks, our results show that it is better to use the CG estimator when (a) the dataset has a significant amount of censoring, approx. 20-25% or more, and (b) the dataset has an adequate size ($\geq$5,000 instances, like the MIMIC-IV and SEER cohorts).
How Do We Select the Copula? Estimation of the copula is inherently difficult due to the non-identifiability of competing risks data [Tsiatis, 1975], however, there exist several strategies to approximate the copula beside the likelihood-approach that we adopt: (a) cross-validation using Harrell’s CI [Emura and Chen, 2016], (b) simulation based on causal inference [Parikh et al., 2022, Mahajan et al., 2024], or (c) using domain knowledge to select the copula a priori [Foomani et al., 2023]. In (a), we first divide $N$ instances into $K$ groups of approximately equal sample sizes, and for each instance $i\in K$, we define a prognostic index, $PI_{i}(C_{\theta},\bm{x}_{i},W)$, which predicts the outcome for $i$, where $W$ are the model parameters. Using cross-validation, we pick the copula $\hat{\theta}$ which maximizes Harrell’s CI over possible values of Kendall’s $\tau$. However, this approach is computationally expensive. In (b), we generate counterfactual data with characteristics similar to the observed data and evaluate the model performance on the simulated data using counterfactual metrics (such as the Survival-$\ell_{1}$ [Foomani et al., 2023]). This approach is less computationally demanding, but requires adequate data and careful hyperparameter tuning [Parikh et al., 2022]. Finally, we can (c) plot the event and censoring distributions to visually inspect the curves. For example, when the censoring survival curve declines faster than the event curve, the event quantile exceeds the censoring quantile at a given time, indicating that a stronger dependence between the event and censoring times increases the likelihood of censoring before the event.
6 Conclusion
In this paper, we have proposed three copula-based metrics for evaluating survival models under dependent censoring. We empirically compared them to existing, independent variants in both synthetic and semi-synthetic datasets. We found that in datasets with adequate data and censoring, the dependent metrics can give error estimates that are closer to the true error, mainly in terms of predictive accuracy, i.e., Brier Score and MAE. We anticipate that others will be able to use these metrics and improve upon our framework.
Acknowledgements.
This research received support from the Natural Science and Engineering Research Council of Canada (NSERC), the Canadian Institute for Advanced Research (CIFAR), and the Alberta Machine Intelligence Institute (Amii).
Appendix A Overview of the appendices
Appendix B provides an overview of the notation. Appendix C describes the main censoring assumptions. Appendix D describes the theory behind copulas. Appendix E describes the existing survival metrics. Appendix F gives details on data generation, preprocessing, survival learners and implementation. Appendix G contains the complete results for the datasets. Appendix H provides derivations of used estimators.
Appendix B Notation
Table 2 provides a list of notations used in this work.
Symbol/Abbr. Definition
$c^{(i)}$ Censoring time of instance $i$
$e^{(i)}$ Event time of instance $i$
$t^{(i)}$ Observed time of instance $i$
$\hat{t}^{(i)}$ Predicted time of instance $i$
$\delta^{(i)}_{k}$ Event bit/indicator, $\delta^{(i)}=1_{e^{(i)}>c^{(i)}}$
$f_{T}$ Probability density function, representing $\text{Pr}(T=t)$
$F_{T}$ Cumulative density function, representing $\text{Pr}(T<t)$
$S_{T}$ Survival function, $S:\mathbb{R}\rightarrow[0,1]$
$\mathcal{D}$ Raw dataset
$\mathcal{D^{\prime}}$ Raw dataset with only uncensored subjects
$\mathcal{D^{\prime\prime}}$ Semi-synthetic dataset with synthetic censoring on $\mathcal{D^{\prime}}$
$\tilde{\mathcal{D}}$ Raw dataset with flipped censor bit
$N$ Number of observations in the dataset
$d$ Number of features/covariates in the dataset
$C$ Censoring rate (%) in the dataset
$\bm{x}^{(i)}$ Covariates of instance $i$
$G(t)$ Censoring distribution
$G_{CoxPH}(t\mid\bm{X})$ Feature-dependent censoring distribution, estimated using CoxPH
$h_{0}(t)$ Baseline hazard function
$h(\cdot\mid\bm{x}^{(i)})$ Hazard function given the covariates $\bm{x}^{(i)}$
$\mathcal{L}$ Objective/loss function
$C_{\theta}$ Copula parameterized by $\theta$
$u_{1},u_{2}$ Inputs to a copula function
Table 2: Table of notation.
Appendix C Censoring assumptions
Random censoring. We can define random censoring as follows: “Subjects who are censored at time t should be representative of all the study subjects who remained at risk at time t with respect to their survival experience.” [Kleinbaum and Klein, 2012, Ch. 1]. In other words, we assume that the death rate (or failure rate) of censored subjects is the same as that of uncensored subjects who remain in the risk set. Random censoring is a stronger assumption and more restrictive than what is called independent censoring, which states that censoring is independent within a subgroup.
Definition 5
Random censoring
$h_{\text{CE}}(t)=h_{\text{NCE}}(t)$ (C.1)
where $h_{\text{CE}}$ is the hazard rate for the censored subjects, and $h_{\text{NCE}}$ is the hazard rate for the uncensored subjects.
Independent censoring. In the absence of competing risks, we can define independent censoring as follows: “Within any subgroup of interest, subjects who are censored at time $t$ should be representative of all subjects in that subgroup who remained at risk at time $t$ with respect to their survival experience” [Kleinbaum and Klein, 2012, Ch. 1]. In other words, censoring is independent provided that it is random within any subgroup given the covariates.
Definition 6
Independent censoring
$h_{A,\text{CE}}(t)$ $=h_{A,\text{NCE}}(t)$ (C.2)
$h_{B,\text{CE}}(t)$ $=h_{B,\text{NCE}}(t)$
where $A$ and $B$ are subgroups, e.g., patients given treatment A and patients given treatment B.
The idea behind random and independent censoring is that, even though the censored subjects were not randomly selected, their survival experience would be expected to be the same as if they had been randomly selected from the risk set at time $t$. Consider a clinical trial studying the recovery time of patients after surgery. The patients’ recovery times are measured, but some individuals leave the study early for reasons unrelated to their recovery, such as relocating to another city, personal commitments, or withdrawal for reasons not related to their health (e.g., family obligations). These individuals are censored, but their reason for leaving the study has nothing to do with the outcome (recovery time). Many of the analytic techniques in survival analysis, e.g., the Kaplan-Meier [Kaplan and Meier, 1958] and Nelson-Aalen [Nelson, 1969, Aalen, 1978] population estimators, and the semiparametric Cox Proportional Hazards (CoxPH) model [Cox, 1972], rely on the assumption of independent censoring for valid inference in the presence of right-censored data [Kleinbaum and Klein, 2012, Ch. 1]. When censoring is not independent, however, this can introduce bias in survival analysis. Bias arises when censored individuals have a higher likelihood of experiencing the event compared to those who are not censored. Consequently, if a significant portion of censored individuals actually had the event (e.g., died), the estimated survival probability at any time $t$ may overestimate the true survival probability at that time. By example of Kleinbaum and Klein [2012, Ch. 1], consider a drug study where some patients are censored due to the occurrence of side effects. The unobserved survival outcomes of those censored because of side effects may not be representative of those who remain in the study. If patients who experience side effects are more susceptible to the health outcome, e.g., death, assuming independent censoring could lead to an overestimation of their survival rate.
Non-informative censoring. Censoring can be non-informative if the distribution of survival times ($T_{E}$) provides no information about the distribution of censoring times ($T_{C}$), and vice versa. Otherwise, the censoring is informative. However, we would still need to know which subjects are censored or not censored. The assumption of non-informative censoring is often justifiable when censoring is independent and/or random [Kleinbaum and Klein, 2012, Ch. 1], but cases exist where the censoring is informative but also random and independent: Consider a cohort study, where each time a patient experiences an event (e.g., death), another patient in the study is randomly selected to be censored. This could be a family member who decides to leave the study after an event occurs. If those who are censored are representative of those who remain in the risk set, then the censoring would be random and independent. However, in this case, the censoring would also be informative, as the censoring mechanism would be linked to the time-to-event distribution (since events trigger censoring). In fact, if this were the only mechanism by which individuals were censored, the distribution of survival times would entirely determine the distribution of censoring times, making it highly informative.
Appendix D Copulas and Definitions
We can view bivariate survival analysis as a competing risks problem, where the occurrence of an event (e.g., death) and censoring (e.g., dropout) are mutually-exclusive incidents. Tsiatis [1975] called this the identifiability issue; competing risks data do not allow one to observe two event times simultaneously (one event censors the other), and hence, the data may not identify the dependence structure between event times. Zheng and Klein [1995] gave a partial solution to the identifiability problem of dependent censoring by an assumed copula between two event times. A copula is a function that links two or more random variables by specifying their dependence structure [Emura and Chen, 2018, Ch. 1]. A mathematician, Abe Sklar, first used the word copula in his study of probabilistic metric space [Sklar, 1959]. In his paper, he gave a mathematical definition of copulas and established the most fundamental theorem about copulas, known as Sklar’s theorem [Sklar, 1959], and demonstrated that any joint cumulative density can be written in terms of a copula over the quantiles of its marginal cumulative densities. Among the various types of copulas, the Archimedean copula is the most used, and enables the modeling of dependencies in arbitrarily high dimensions using a single parameter. This family of copulas is defined by a convex generator function, $\varphi:[0,1]\to[0,\infty]$, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [Emura and Chen, 2018, Ch. 3]. Given $T$ and $C$, we can define an Archimedean copula as:
$\begin{split}P(T>t,C>c)&=\varphi^{-1}\big{[}\varphi\{S_{T}(t)\}\\ &\quad+\varphi\{S_{C}(c)\}\big{]},\end{split}$ (D.1)
where $\varphi$ is the generator function, and $S_{T}(t)=P(T>t)$ and $S_{C}(c)=P(C>c)$ are the marginals survival functions. In the bivariate case, $\varphi$ describes the dependency structure (i.e., correlation) between two random variables, e.g.$T_{E}$ and $T_{C}$, and the (rank) correlation can described with Kendall’s tau ($\tau$):
$\begin{split}\tau=&\quad\text{Pr}\{\ (T_{2}-T_{1})(C_{2}-C_{1})\geq 0\}\\ &-\text{Pr}\{\ (T_{2}-T_{1})(C_{2}-C_{1})<0\},\end{split}$ (D.2)
where ($T_{1}$, $C_{1}$) and ($T_{2}$, $C_{2}$) are sampled from a (survival) copula. The Clayton [Clayton, 1978] and Frank [Frank, 1979] copulas are two common Archimedean copulas in survival literature. Within these families, the copula $C_{\theta}$ is parameterized by $\theta$, interpreted as the degree of dependence between $T_{E}$ and $T_{C}$ in the bivariate case. When $\theta\approx 0$, it yields the independence copula, thus a larger value of $\theta$ implies greater dependency between the marginal distributions. Formally, we write $C(u_{1},\ldots,u_{d}):[0,1]^{d}\rightarrow[0,1]$, which is a $d$-dimensional copula (i.e., where $d$ is the number of variables) with $u_{1},\ldots,u_{d}$ uniform marginal probabilities and the following properties Nelsen [2006]:
1. Groundedness: If there exists an $i\in\{1,\dots,m\}$ such that $u_{i}=0$, then
$C(u_{1},\dots,u_{m})=0.$
2. Uniform Margins: For all $i\in\{1,\dots,m\}$, if $\forall j\neq i\Rightarrow u_{j}=1$, then
$C(u_{1},\dots,u_{m})=u_{i}.$
3. $d$-Increasingness: For all $u=(u_{1},\dots,u_{m})$ and $v=(v_{1},\dots,v_{m})$ where $u_{i}<v_{i}$ for all $i=1,\dots,m$, the following holds:
$\sum_{\ell\in\{0,1\}^{m}}(-1)^{\ell_{1}+\dots+\ell_{m}}C\left(u_{1}^{\ell_{1}}v_{1}^{1-\ell_{1}},\dots,u_{m}^{\ell_{m}}v_{m}^{1-\ell_{m}}\right)\geq 0.$
Let $T_{E}$ be the event time, $T_{C}$ is the censoring time, $\bm{x}$ is a vector of covariates, and $S_{T_{E}}(t\mid\bm{x})=\text{Pr}(T_{E}>t_{e}\mid\bm{x})$ and $S_{T_{C}}(u\mid\bm{x})=\text{Pr}(T_{C}>t_{c}\mid\bm{x})$ are the marginal survival functions given $\bm{x}$. We can then define a bivariate (survival) copula $C_{\theta}$ [Nelsen, 2006] that describes the degree of dependence between $T_{E}$ and $T_{C}$ as follows:
$\text{Pr}(T_{E}>t_{e},T_{C}>t_{c}\mid\mathbf{x})\ =\ C_{\theta}\{\,S_{T_{E}}(t_{e}\mid\bm{x}),\,S_{T_{C}}(t_{c}\mid\bm{x})\,\}.$ (D.3)
Kendall’s tau ($\tau$) is a well-known measure to assess the dependence between $T_{E}$ and $T_{C}$, which is defined by [Emura and Chen, 2018, Ch. 3]:
$\begin{split}\tau=&\quad\text{Pr}\{\ (T_{2}-T_{1})(C_{2}-C_{1})\geq 0\mid\bm{x}\ \}-\text{Pr}\{\ (T_{2}-T_{1})(C_{2}-C_{1})<0\mid\bm{x}\ \},\end{split}$ (D.4)
where ($T_{1}$, $C_{1}$) and ($T_{2}$, $C_{2}$) are sampled from the model (D.3). Kendall’s $\tau$ can be solely expressed as a function of the copula $C_{\theta}$ by [Emura and Chen, 2018, Ch. 3]:
$\tau_{\theta}=4\int_{0}^{1}\int_{0}^{1}C_{\theta}(u,v)\,C_{\theta}(du,dv),$ (D.5)
which implies that Kendall’s $\tau$ does not depend on the marginal survival functions. An Archimedean copula is a type of copula that is defined using a generator function, which allows for a flexible way to model dependence structures between random variables, e.g.$T_{E}$ and $T_{C}$. Such a copula is defined as [Emura and Chen, 2018, Ch. 3]:
$C_{\theta}(u,v)\ =\ \varphi_{\theta}^{-1}\left(\,\varphi_{\theta}(u)+\varphi_{\theta}(v)\,\right),$ (D.6)
where the function $\varphi_{\theta}:[0,1]\rightarrow[0,\infty]$ is called a generator of the copula, which is continuous and strictly decreasing from $\varphi_{\theta}(0)\geq 0$ to $\varphi_{\theta}(1)=0$. The following Archimedean copulas are considered in this work:
  • The Independence copula: $C(u,v)=uv$.
  • The Clayton copula [Clayton, 1978]:
    $C_{\theta}(u,v)=\left(u^{-\theta}+v^{-\theta}-1\right)^{-1/\theta},\quad\theta\geq 0.$
  • The Frank copula [Frank, 1979]:
    $C_{\theta}(u,v)=-\frac{1}{\theta}\log\left[1+\frac{(e^{-\theta u}-1)(e^{-\theta v}-1)}{e^{-\theta}-1}\right].$
    $\quad\theta\neq 0.$
Appendix E Existing Evaluation Metrics
E.1 Concordance Index
The Harrell’s concordance index (CI) quantifies the proportion of concordant pairs among all comparable pairs [Harrell Jr et al., 1996]. Given an estimated risk function $\hat{\eta}(\cdot)$ and corresponding labels, CI is computed as:
$\text{CI}(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{\eta}(\cdot))\ =\ \frac{\sum_{i,j:i\neq j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j:i\neq j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]}.$
Some definitions of CI [Hartman et al., 2023] incorporate tie-breaking rules for cases where risk scores are identical. Specifically, a score of 0.5 may be assigned to tied pairs in the numerator. However, for simplicity, we omit ties in our discussion.
As shown in Uno et al. [2011], this calculation converges to a censoring-dependent probability:
$P(\ \hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})\mid t_{i}<t_{j},t_{i}<\min(t_{e,j},t_{c,j})\ ),$
which differs from the intended probability:
$P(\ \hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})\mid t_{i}<t_{j}\ ),$
To address this bias, Uno et al. [2011] proposed the IPCW-weighted CI, computed as [^1] The proposed IPCW-CI also comes with a truncation time $\tau$ as a parameter – which means we only focus on the accuracy of ranking within $[0,\tau]$. Here, for consistency with other CIs, we do not consider such a truncation time.:
$\text{CI}_{\text{IPCW}}(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{\eta}(\cdot))\ =\ \frac{\sum_{i,j:i\neq j}\delta_{i}\cdot\{\hat{G}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j:i\neq j}\delta_{i}\cdot\{\hat{G}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]},$ (E.1)
where $\hat{G}(t)$ represents the marginal survival probability for censoring at time $t$, typically estimated using the Kaplan-Meier (KM) estimator. The inverse probability censoring weight (IPCW) is given by $1/\hat{G}(t)$. In cases where censoring probability depends on features (i.e., $\hat{G}(t\mid\bm{x}_{i})$), a survival model such as CoxPH can be used for estimation. These two metrics are implemented in package SurvivalEVAL [Qi et al., 2023b].
E.2 Integrated Brier Score
The Brier Score (BS) quantifies the accuracy of probabilistic survival predictions at a specific time $t^{*}$ [Graf et al., 1999]. It measures the squared error between the predicted survival probability and the observed survival status at $t^{*}$:
To incorporate the censored subjects in the calculation, BS also requires estimating the IPCW to uniformly transfer a censored subject’s weight (originally 1) to subjects with known status at that time [Vock et al., 2016]. Formally:
$\text{BS}\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{S}(\cdot),\hat{G}(\cdot),t^{*}\right)\ =\ \frac{1}{N}\sum_{i}\left(\frac{\delta_{i}\cdot\mathbbm{1}[t_{i}\leq t^{*}]\cdot\hat{S}(t^{*}\mid\bm{x}_{i})^{2}}{\hat{G}(t_{i})}+\frac{\mathbbm{1}[t_{i}>t^{*}]\cdot(1-\hat{S}(t^{*}\mid\bm{x}_{i}))^{2}}{G(t^{*})}\right),$
The integrated Brier score (IBS) is the expectation of BS over all time, defined as:
[…]
where $t_{\text{max}}$ is usually the maximum event time of the combined training and validation datasets. This metric is implemented in package SurvivalEVAL [Qi et al., 2023b].
E.3 Mean Absolute Error
Survival prediction shares similarities with regression, as it estimates a continuous event time given a feature set $\bm{x}_{i}$. Therefore, mean absolute error (MAE) serves as a useful evaluation metric.
MAE-Margin [Haider et al., 2020] estimates a “best guess” event time for censored subjects using the KM estimator [Kaplan and Meier, 1958]. For a subject censored at time $t_{i}$, the margin event time is computed as:
[…]
where $S_{\text{KM}}(t)$ is the KM estimation that is typically derived from the training dataset.
Additionally, for each censored subject, Haider et al. [2020] suggested we use a confidence weight $\omega_{i}=1-S_{\text{KM}}(t_{c,i})$ for the error calculated based on the margin value. This weight $\omega_{i}$ yields lower confidence for early censoring subjects and higher confidence for late censoring data. Of course, we set the weights for uncensored subjects $i$ to $w_{i}=1$ as we have full confidence in those error calculations. The overall MAE margin after a re-weighting scheme is:
$\text{MAE}_{\text{margin}}\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{t}_{i}\right)=\frac{1}{\sum_{i=1}^{N}\omega_{i}}\sum_{i=1}^{N}\omega_{i}\left|[(1-\delta_{i})\cdot t_{e}^{\text{margin}}(t_{i})+\delta_{i}\cdot t_{i}]-\hat{t}_{i}\right|\ .$
This metric is implemented in package SurvivalEVAL [Qi et al., 2023b].
Appendix F Experimental Details
F.1 Data generation and preprocessing
Dataset #Instances #Features %Censored #Event Max Event $t$
METABRIC 1,902 9 (9) 42.1% 1102 355
[…]
MIMIC-IV (all causes) 38,520 91 (103) 66.7% 12845 4404
[…]
SEER (liver) 82,841 14 (14) 37.6% 51704 225
SEER (stomach) 100,360 14 (14) 43.4% 56807 226
Table 3: Overview of the raw datasets. The datasets are sorted by the number of instances. The number of features before performing one-hot encoding, with brackets (the number of features after one-hot encoding).
In this section, we describe how the synthetic dataset is generated and how we preprocessed the raw survival datasets. Table 3 shows an overview of the raw datasets.
Synthetic: This paper utilizes a linear synthetic dataset with two competing events (i.e., event and censoring), both based on the semi-parametric Weibull distribution as discussed in Foomani et al. [2023]. The datasets are characterized by the conditional hazard function and conditional survival probability of an event as follows:
$h_{T\mid X}(t\mid X)$ $=\left(\frac{v}{\rho}\right)\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(X)\right),$ (F.1)
$S_{T\mid X}(t\mid X)$ $=\exp\left(-\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(X)\right)\right),$ (F.2)
where $v$ and $\rho$ represent the shape and scale parameters of the Weibull distribution, respectively. We use $v_{E}=4$, $v_{C}=6$, $\rho_{E}=17$, $\rho_{C}=19$. $g_{\Psi}(X)$ represents the risk function, which is a linear function in this work. The data generation process begins by initializing a copula with a predefined parameter $\theta$. We then sample $N$ sets of CDF observations from the copula, each containing two CDF probabilities (corresponding to each competing event). From these probabilities, we then calculate the corresponding times using the inverse function of (F.2). The data generation process is described in Algorithm 2 in Foomani et al. [2023].
METABRIC: The Molecular Taxonomy of Breast Cancer International Consortium (METABRIC) dataset [Curtis et al., 2012] contains survival information for breast cancer patients. This dataset includes a diverse range of feature sets that encompass clinical traits, expression profiles, copy number variation (CNV) profiles, and single nucleotide polymorphism (SNP) genotypes. All these features are derived from breast tumor samples collected during the METABRIC trial. The dataset can be downloaded from (https://www.cbioportal.org/study/summary?id=brca_metabric), and it does not have any missing values.
MIMIC-IV: The Medical Information Mart for Intensive Care (MIMIC)-IV [Johnson et al., 2023] contains critical care data from patients admitted to hospital and intensive care units (ICU). We create a version using the MIMIC-IV database which contains patients that are alive at least 24 hours after being admitted to ICU. Their date of death is derived from hospital records or state records, which means, the cause of mortality is not limited to the reason for ICU admission. We follow the instructions in [Qi et al., 2023a] to process the dataset and the code is available in their GitHub repository: https://github.com/shi-ang/CensoredMAE.
SEER: The SEER Program dataset [LAG et al., 2007] is a comprehensive collection of cancer patient data from approximately 49% of the U.S. population. Our study focuses on three subsets: SEER-brain, SEER-liver, and SEER-stomach, containing data on patients diagnosed with brain, liver, and stomach cancers, respectively. The goal is to model the time from diagnosis to failure events, such as death or disease progression, after preprocessing steps outlined by [Qi et al., 2024]. Features analyzed include age, sex, cancer stage, grade, and various clinical factors. We excluded features with over 70% missing data, patients without follow-up times, duplicates, and those with a survival time of zero. The SEER cohort is available for download at www.seer.cancer.gov.
F.2 Survival learners
In this section, we will describe the implementation and hyperparameters of the survival learners.
CoxPH: The Cox Proportional Hazards model (CoxPH) is a semi-parametric survival analysis model [Cox, 1972], consisting of a non-parametric baseline hazard function $h_{0}(t)$ and a parametric partial hazard function $f(\bm{\theta},\bm{x}_{i})$, typically estimated using the Breslow method [Breslow, 1975]. The hazard function is $h(t|\bm{x}_{i})=h_{0}(t)\exp(f(\bm{\theta},\bm{x}_{i}))$, where $f$ is linear. The MLE of $\hat{\bm{\theta}}$ maximizes the Cox partial log-likelihood [Cox, 1972].
GBSA: GBSA [Ridgeway, 1999] is a likelihood-based boosting model that optimizes an $\ell_{2}$-norm penalized Cox partial log-likelihood [Cox, 1972]. It updates candidate variables flexibly at each boosting step using an offset-based gradient boosting method, unlike traditional gradient boosting, which updates either a single component or all features for fitting the gradient.
RSF: Random Survival Forests (RSF) [Ishwaran et al., 2008] extend decision trees for survival analysis by constructing multiple trees on bootstrapped samples. Each tree predicts an individual’s survival function, with splits based on criteria like the log-rank test. RSF handles censored data as right-censored and accounts for it during tree construction, with final predictions as the average of all trees’ survival functions.
DeepSurv: DeepSurv [Katzman et al., 2018] is a neural network implementation of the CoxPH model of the form $h\left(t|\bm{x}_{i}\right)=h_{0}\left(t\right)\exp\left(f\left(\bm{\theta},\bm{x}_{i}\right)\right)$, where $f\left(\bm{\theta},\bm{x}_{i}\right)$ denotes a risk score as a nonlinear function of the features, i.e.$f\left(\bm{\theta},\bm{x}_{i}\right)=\sigma(\bm{x}_{i}\bm{\theta})$, where $\sigma$ is a nonlinear activation function. Popular functions include the hyperbolic tangent (Tanh) function or the Rectified Linear Unit (ReLU). The MLE for $\hat{\bm{\theta}}$ is derived by numerically maximizing the partial Cox log-likelihood.
MTLR: Multi-Task Logistic Regression (MTLR) Yu et al. [2011] is a neural network implementation of a discrete-time model that predicts the survival distribution using a sequence of dependent logistic regressors. The number of discrete times is determined by the square root of the number of uncensored patients, and we use quantiles to divide those uncensored instances evenly into each time interval.
Table 4 reports the selected hyperparameters for the survival learners.
Model Parameter Value
CoxPH alpha 0.0001
ties ’Breslow’
n_iter 100
tol 1e-9
GBSA n_estimators 100
max_depth 1
min_samples_split 2
min_samples_leaf 1
max_features ’sqrt’
sub_sample 0.8
random_state 0
RSF n_estimators 100
max_depth 1
min_samples_split 2
min_samples_leaf 1
max_features ’sqrt’
random_state 0
DeepSurv hidden_size 100
verbose False
lr 0.001
c1 0.01
num_epochs 1000
batch_size 32
dropout 0.25
early_stop True
patience 10
MTLR verbose False
lr 0.001
c1 0.01
num_epochs 1000
batch_size 32
early_stop True
patience 10
Table 4: Hyperparameters for the survival learners.
F.3 Implementation Details
Copula model. To validate our approach in semi-synthetic datasets, we train two copula models, Clayton and Frank, alongside two nonlinear Weibull proportional-hazards models for the marginal distributions $T_{E}$ and $T_{C}$, respectively. The Weibull models use a multilayer perceptron (MLP) as the backbone network architecture. We use a single layer with 32 hidden nodes and the ReLU activation function. For all experiments, we train the models using the Adam optimizer [Kingma and Ba, 2017] with a learning rate of 1e-3, batch normalization enabled, and a batch size of 1024.
Loss function. We estimate the model parameters using maximum likelihood estimation (MLE) under dependent censoring, as described in Foomani et al. [2023]. Given survival data $\mathcal{D}$, let $f_{T\mid X}(\cdot)$ and $F_{T\mid X}(\cdot)$ represent the conditional density and cumulative distribution functions, respectively, over the event horizon, with the event distribution $T_{E}$ and the censoring distribution $T_{C}$. Under independence censoring, the likelihood of the $i$-th instance experiencing the event can then be determined:
$\mathcal{L}(\mathcal{D})=\prod_{i=1}^{N}\left[\text{Pr}(T_{E}=t_{i},T_{C}\geq t_{i}\mid\bm{x}^{(i)})^{\delta^{(i)}}\times\text{Pr}(T_{E}\geq t_{i},T_{C}=t_{i}\mid\bm{x}^{(i)})^{1-\delta^{(i)}}\right]$ (F.3)
However, when $T_{E}$ and $T_{C}$ are no longer independent, we can no longer rely on this clean decomposition of the likelihood. Instead, we introduce a copula $C(u_{1},u_{2})$ to Equation F.3 to adjust for dependent censoring by applying Lemma 2 from Foomani et al. [2023]. Now, as a function of $C(u_{1},u_{2})$ and taking the sum of logs, the (dependent) likelihood has the following form:
$\mathcal{L}(\mathcal{D})=\sum_{i=1}^{N}\delta^{(i)}\;\text{log}\left[f_{T_{E}|\bm{x}^{(i)}}(t_{i}|\bm{x}^{(i)})\right]\;+\;\delta^{(i)}\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T_{E}\rvert\bm{x}^{(i)}}(t_{i}\rvert\bm{x}^{(i)})}_{u_{2}=S_{T_{C}\rvert\bm{x}^{(i)}}(t_{i}\rvert\bm{x}^{(i)})}\right]$ (F.4)
$+\;(1-\delta^{(i)})\;\text{log}\left[f_{T_{C}|\bm{x}^{(i)}}(t_{i}|\bm{x}^{(i)})\right]+(1-\delta^{(i)})\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T_{E}\rvert\bm{x}^{(i)}}(t_{i}\rvert\bm{x}^{(i)})}_{u_{2}=S_{T_{C}\rvert\bm{x}^{(i)}}(t_{i}\rvert\bm{x}^{(i)})}\right]$
In this expression, the first term represents the log-likelihood of observing the event at time $t_{i}$. The second term captures the conditional probability of observing the censoring time after the event time, given that the event occurred at $t_{i}$. The third and fourth terms, by symmetry, represent the same quantities for the censoring time. Given $C(u_{1},u_{2})$ is an Archimedean copula (e.g., Clayton or Frank), the partial derivatives in Equation F.4 have closed-form solutions, allowing the log-likelihood to be expressed in a closed form that can be optimized using gradient-based methods. To estimate F.4, we adopt the approach as described in Algorithm 1 in Foomani et al. [2023].
Copula selection. After training, for each model (marginals and copula), we choose the model that minimizes the Akaike Information Criterion (AIC) given the model’s negative log-likelihood (NLL) and its number of parameters. The AIC is a measure used to compare models, taking into account both the goodness of fit (via the likelihood) and model complexity (via the number of parameters). This offers a practical, data-driven approach to copula selection, as minimizing the NLL provides an objective criterion for which model best captures the dependency structure explicitly [Emura and Chen, 2018, Ch. 4]. We then extract $C_{\theta}$ from the chosen copula model and use that to calculate the dependent evaluation metrics.
Reproducibility. All experiments were implemented and executed in Python 3.9 with PyTorch 1.13.1, NumPy 1.24.3, and Pandas 1.5.3. The Copula-Graphic estimator was implemented using the “compound.Cox” version 3.32 package. Tensor computations were carried out in double precision (fp64). The experiments were conducted on a single workstation equipped with an Intel Core i9-10980XE 3.00GHz CPU, 64GB of RAM, and an NVIDIA GeForce RTX 3090 GPU. Model training was performed on the GPU using CUDA 11.7. All datasets utilized in this study are publicly available. Detailed instructions for reproducing the results are included in the supplementary code.
Appendix G Complete results
G.1 Synthetic data
Figures 5-7 demonstrate the bias, i.e., the proximity to the true error, of established baseline metrics and the proposed dependent metrics as a function of the dependence (Kendall’s $\tau$) under a known copula. We plot the mean biases across 10 experiments with different random seeds. Data is generated from a DGP with sample sizes $N=\{1000,5000,10000\}$ and $d=10$. Across all metrics, we generally see larger errors with high dependence ($\tau$). Concerning the small dataset ($N=1,000$), the CI-Dep variant exhibits higher error than Harrell’s CI for both copulas under low to medium dependency ($\tau\leq 0.5)$, however, the CI-Dep improves on both Harrell’s and Uno’s under high dependency ($\tau>0.5)$). The same is the case for the IBS and MAE results, especially under low dependency. In the medium dataset ($N=5,000$), the dependent metrics consistently outperform their independent counterparts across all levels of $\tau$. This trend becomes even more pronounced in the large dataset ($N=10,000$), where the dependent metrics show a significant improvement over the independent ones. These results highlight the increasing impact of dependent censoring as dataset size grows, further demonstrating the limitations of the independent censoring assumption in larger datasets.
Figure 5: Plot of the bias of established baseline metrics and the proposed dependent metrics. $N=1000$, $d=10$.
Figure 6: Plot of the bias of established baseline metrics and the proposed dependent metrics. $N=5000$, $d=10$.
Figure 7: Plot of the bias of established baseline metrics and the proposed dependent metrics. $N=10000$, $d=10$.
G.2 Semi-synthetic data
Table 5 shows the results in the 6 semi-synthetic datasets. We see that CI-Dep can improve on Uno’s CI (i.e., give lower bias) in cases with large sample size, e.g., the SEER datasets, but Harrell’s CI is generally the better choice among the two. Concerning the IBS results, the IBS-Dep offers a solid alternative to the IPCW method, as it gives notable improvements in several datasets. Concerning the MAE results, the MAE-Dep is aligned with the MAE-Margin in the smaller datasets, i.e., METABRIC and MIMIC-IV, but gives notable improvements in the larger SEER datasets.
Figures 8-13 shows the ranking accuracy, i.e., the number of times the resp. metric correctly identified the ordering of the top-3 survival learners by the true metric, over 5 experiments in 6 datasets. We see general alignment between the metrics, but IBS-Dep and MAE-Dep give notable improvements over their independent variants in the SEER (liver) dataset.
Dataset Strategy Discrimination (CI) Accuracy (IBS) Accuracy (MAE)
Harrell’s Uno’s Dependent IPCW Dependent Uncens. Hinge Pseudo Margin Dependent
METABRIC ($N$=1,102, $C$=3.8%) Original 0.10$\pm$0.08 0.39$\pm$0.14 0.48$\pm$0.25 1.23$\pm$0.85 1.17$\pm$0.85 3.24$\pm$0.55 1.40$\pm$0.21 0.25$\pm$0.21 0.31$\pm$0.23 0.33$\pm$0.22
Top 5 0.11$\pm$0.08 0.51$\pm$0.17 0.58$\pm$0.23 1.22$\pm$0.85 1.17$\pm$0.85 3.21$\pm$0.55 1.40$\pm$0.21 0.25$\pm$0.20 0.31$\pm$0.23 0.30$\pm$0.23
Top 10 0.10$\pm$0.08 0.43$\pm$0.11 0.49$\pm$0.16 1.23$\pm$0.85 1.17$\pm$0.85 3.26$\pm$0.54 1.40$\pm$0.21 0.25$\pm$0.21 0.32$\pm$0.23 0.31$\pm$0.23
Rand. 25% 0.10$\pm$0.08 2.34$\pm$0.68 2.31$\pm$0.70 1.25$\pm$0.88 1.18$\pm$0.88 3.45$\pm$0.48 1.40$\pm$0.21 0.27$\pm$0.22 0.34$\pm$0.23 0.33$\pm$0.23
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original 0.65$\pm$0.21 0.66$\pm$0.21 0.66$\pm$0.21 0.76$\pm$0.17 0.92$\pm$0.17 36.73$\pm$4.98 62.28$\pm$5.12 37.18$\pm$4.98 37.18$\pm$4.98 37.34$\pm$4.96
Top 5 1.11$\pm$0.19 1.60$\pm$0.23 1.60$\pm$0.23 0.82$\pm$0.19 1.12$\pm$0.20 39.35$\pm$5.88 65.63$\pm$5.97 39.56$\pm$5.87 39.56$\pm$5.87 39.72$\pm$5.87
Top 10 1.02$\pm$0.17 1.29$\pm$0.21 1.29$\pm$0.21 0.79$\pm$0.19 1.02$\pm$0.20 37.71$\pm$5.57 63.72$\pm$5.69 38.13$\pm$5.60 38.13$\pm$5.60 38.33$\pm$5.59
Rand. 25% 0.45$\pm$0.17 0.48$\pm$0.18 0.48$\pm$0.18 0.81$\pm$0.18 1.07$\pm$0.20 36.95$\pm$3.93 62.79$\pm$4.10 37.25$\pm$3.94 37.25$\pm$3.94 37.46$\pm$3.95
MIMIC-IV (hospital) ($N$=6,780, $C$=74.3%) Original 0.74$\pm$0.43 1.27$\pm$0.83 2.10$\pm$1.54 8.75$\pm$1.50 6.57$\pm$1.44 3.48$\pm$0.25 6.44$\pm$0.27 3.11$\pm$0.27 4.54$\pm$0.26 4.36$\pm$0.24
Top 5 0.79$\pm$0.55 1.16$\pm$0.56 3.36$\pm$2.74 8.80$\pm$1.51 6.42$\pm$1.45 3.51$\pm$0.24 6.45$\pm$0.26 3.10$\pm$0.26 4.65$\pm$0.25 4.33$\pm$0.18
Top 10 0.71$\pm$0.47 1.10$\pm$0.65 2.46$\pm$1.69 8.73$\pm$1.50 6.57$\pm$1.44 3.50$\pm$0.27 6.45$\pm$0.27 3.11$\pm$0.27 4.60$\pm$0.28 4.40$\pm$0.25
Rand. 25% 1.44$\pm$0.58 1.77$\pm$1.02 2.70$\pm$1.44 10.38$\pm$0.76 7.91$\pm$0.90 3.57$\pm$0.20 6.51$\pm$0.19 3.06$\pm$0.19 4.79$\pm$0.19 4.62$\pm$0.20
SEER (brain) ($N$=44,137, $C$=9.0%) Original 0.45$\pm$0.08 0.86$\pm$0.10 0.54$\pm$0.11 0.44$\pm$0.14 0.42$\pm$0.34 1.79$\pm$0.06 1.21$\pm$0.04 1.50$\pm$0.06 1.11$\pm$0.04 1.13$\pm$0.04
Top 5 0.25$\pm$0.06 0.87$\pm$0.13 0.63$\pm$0.14 0.44$\pm$0.15 0.37$\pm$0.33 1.78$\pm$0.06 1.22$\pm$0.04 1.50$\pm$0.06 1.11$\pm$0.04 1.13$\pm$0.04
Top 10 0.49$\pm$0.07 0.88$\pm$0.11 0.57$\pm$0.12 0.44$\pm$0.15 0.42$\pm$0.34 1.80$\pm$0.06 1.21$\pm$0.04 1.50$\pm$0.06 1.11$\pm$0.04 1.13$\pm$0.04
Rand. 25% 2.61$\pm$2.38 2.83$\pm$1.72 2.72$\pm$1.75 0.48$\pm$0.12 0.24$\pm$0.09 1.89$\pm$0.09 1.30$\pm$0.10 1.45$\pm$0.11 1.05$\pm$0.10 1.07$\pm$0.10
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.77$\pm$0.10 0.57$\pm$0.14 0.76$\pm$0.13 0.33$\pm$0.15 0.25$\pm$0.11 2.59$\pm$0.07 2.50$\pm$0.04 2.47$\pm$0.07 1.48$\pm$0.06 1.44$\pm$0.05
Top 5 1.86$\pm$0.12 1.70$\pm$0.15 1.76$\pm$0.13 0.28$\pm$0.14 0.47$\pm$0.21 2.80$\pm$0.05 2.69$\pm$0.06 2.25$\pm$0.07 1.25$\pm$0.06 0.85$\pm$0.47
Top 10 0.81$\pm$0.08 0.90$\pm$0.09 1.10$\pm$0.07 0.33$\pm$0.15 0.25$\pm$0.11 2.61$\pm$0.08 2.51$\pm$0.04 2.46$\pm$0.06 1.47$\pm$0.05 1.43$\pm$0.05
Rand. 25% 0.89$\pm$0.52 1.80$\pm$1.23 1.68$\pm$1.21 0.45$\pm$0.21 0.35$\pm$0.17 2.75$\pm$0.11 2.49$\pm$0.08 2.50$\pm$0.14 1.49$\pm$0.09 1.47$\pm$0.11
SEER (stomach) ($N$=56,807, $C$=21.9%) Original 0.43$\pm$0.09 0.30$\pm$0.10 0.49$\pm$0.13 0.27$\pm$0.08 0.22$\pm$0.14 1.21$\pm$0.05 2.12$\pm$0.03 2.19$\pm$0.10 1.24$\pm$0.06 1.15$\pm$0.18
Top 5 0.88$\pm$0.05 0.62$\pm$0.08 0.71$\pm$0.05 0.26$\pm$0.08 0.17$\pm$0.11 1.17$\pm$0.06 2.14$\pm$0.04 2.17$\pm$0.10 1.22$\pm$0.06 1.21$\pm$0.07
Top 10 0.37$\pm$0.09 0.29$\pm$0.08 0.43$\pm$0.14 0.28$\pm$0.08 0.22$\pm$0.17 1.18$\pm$0.06 2.11$\pm$0.04 2.21$\pm$0.10 1.26$\pm$0.06 1.15$\pm$0.18
Rand. 25% 0.86$\pm$0.51 1.47$\pm$1.01 1.37$\pm$0.97 0.27$\pm$0.09 0.16$\pm$0.04 1.13$\pm$0.11 2.12$\pm$0.10 2.20$\pm$0.13 1.28$\pm$0.11 1.25$\pm$0.10
Table 5: Full comparison of established survival metrics and the proposed dependent metrics in terms of the mean ($\pm$ SD.) bias, i.e., the proximity to the true error, averaged over 5 survival learners and 5 experiments. $N$ and $C$ are the number of instances and censoring rate, respectively, after preprocessing. CI and IBS results are multiplied by 100.
Figure 8: Plot the ranking accuracy over 5 experiments in the METABRIC dataset.
Figure 9: Plot the ranking accuracy over 5 experiments in the MIMIC-IV (all) dataset.
Figure 10: Plot the ranking accuracy over 5 experiments in the MIMIC-IV (hospital) dataset.
Figure 11: Plot the ranking accuracy over 5 experiments in the SEER (brain) dataset.
Figure 12: Plot the ranking accuracy over 5 experiments in the SEER (liver) dataset.
Figure 13: Plot the ranking accuracy over 5 experiments in the SEER (stomach) dataset.
Appendix H Derivations
H.1 The Kaplan-Meier estimator
This derivation is reproduced from [Emura and Chen, 2018, Ch. 2].
Let $(t_{i},\delta_{i}),\;i=1,\ldots,n$, be survival data without covariates, where $t_{i}=\min(T_{i},U_{i})$, $\delta_{i}=I(T_{i}\leq U_{i})$, and $I(\cdot)$ is the censoring bit. We can then consider the survival function $S(t)$ as a function that decreases in steps, with discontinuities occurring only at the observed times of death. In this case, the survival function can be expressed as
$S(t)=\text{Pr}(T\geq t)=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(T=t_{i})\right)\text{Pr}(T\geq t_{i})$ (H.1)
Second, suppose that $T$ and $C$ are independent. Then, we can write
$S(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(T=t_{i},C\geq t_{i})\right)\text{Pr}(T\geq t_{i},C\geq t_{i})$ (H.2)
$=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(\min(T,C)=t_{i},T\leq C)\right)\text{Pr}(\min(T,C)\geq t_{i})$ (H.3)
Finally, we replace the probability ratio of the last expression by its estimate to obtain
$S(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{\sum_{\ell=1}^{N}I(t_{\ell}=t_{i},\delta_{\ell}=1)/n}{\sum_{\ell=1}^{N}I(t_{\ell}\geq t_{i})/n}\right)$ (H.4)
$=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{1}{n_{i}}\right)$ (H.5)
H.2 The Copula-Graphic estimator
This derivation is reproduced from [Emura and Chen, 2018, Ch. 4].
Consider an Archimedean copula model
$P(T>t,C>c)=\varphi^{-1}\big{[}\varphi\{S_{T}(t)\}+\varphi\{S_{C}(c)\}\big{]},$ (H.6)
where $\phi_{\theta}:[0,1]^{d}\rightarrow[0,\infty]$ is a generator function, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [Emura and Chen, 2018, Ch. 3], and $S_{T}(t)=\text{Pr}(T>t)$ and $S_{C}(t)=\text{Pr}(C>c)$ are the marginal survival functions. Let $(t_{i},\delta_{i}),\;i=1,\ldots,n$, be survival data without covariates, where $t_{i}=\min(T_{i},U_{i})$, $\delta_{i}=I(T_{i}\leq U_{i})$, and $I(\cdot)$ is the censoring bit. Assume that $S_{T}(t)$ is a decreasing step function with jumps at event times. Thus, $\delta_{i}=1$ implies $S_{T}(t_{i})\neq S_{T}(t_{i}-dt)$ and $S_{U}(t_{i})=S_{U}(t_{i}-dt)$. Setting $t=u=t_{i}$ in Equation H.6, we have
$\text{Pr}(T\geq t_{i},C\geq t_{i})=\text{Pr}(S_{T}(t_{i}))+\text{Pr}(S_{C}(t_{i}))$ (H.7)
In the left-side of the preceding equation, we estimate $\text{Pr}(T>t_{i},C>t_{i})$ by ($n_{i}-1)/n$, where $n_{i}-1=\sum_{\ell=1}^{n}I(t_{\ell}>t_{i})$ is the number of event-free instances at time $t_{i}$. Then we have
$\phi_{\theta}(\frac{n_{i}-1}{n})=\phi_{\theta}\{S_{T}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i})\}.$ (H.8)
We set $t=c=t_{i}-dt$ in Equation H.6 and then estimate $\text{Pr}(T>t_{i}-dt,C>t_{i}-dt)$ by $n_{i}/n$. Then,
$\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{T}(t_{i}-dt)\}+\phi_{\theta}\{S_{C}(t_{i})\},\delta_{i}=1$ (H.9)
Equations H.8 and H.9 result in the system of difference equations
$\phi_{\theta}(\frac{n_{i}-1}{n})-\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{T}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i}-dt)\},\delta_{i}=1$ (H.10)
We apply the standard constraint that $S_{T}(t_{i}-dt)=1$ when $t_{i}$ is the earliest event time. Under this condition, the solution to the differential equations is
$\phi_{\theta}\{S_{T}(t)\}$ $=\sum_{t_{i}\leq t,\delta_{i}=1}[\phi_{\theta}\{S_{T}(t_{i})\}-\phi_{\delta}\{S_{T}(t_{i}-dt)\}]$ (H.11)
$=\sum_{t_{i}\leq t,\delta_{i}=1}\phi_{\delta}\big{(}\frac{n_{i}-1}{n}\big{)}-\phi_{\delta}(\frac{n_{i}}{n})$ (H.12)
which is equivalent to the CG estimator. Under the independence copula, given by $\phi_{\delta}(t)=-log(t)$, the CG estimator is equivalent to the Kaplan-Meier estimator. Substituting for the Clayton copula [Clayton, 1978], the CG estimator is written as
$S_{T}(t)=\Bigl{[}1+\sum_{t_{i}\leq t,\delta_{i}=1}\Bigl{\{}(\frac{n_{i}-1}{n})^{-\theta}-(\frac{n_{i}}{n})^{-\theta}\Bigr{\}}\Bigr{]}^{-1/\theta}$ (H.13)