### v3

# Overcoming Dependent Censoring in the Evaluation of Survival Models

Abstract
Conventional survival metrics, such as Harrell’s concordance index (CI) and the Brier Score, rely on the independent censoring assumption for valid inference with right-censored data. However, in the presence of so-called dependent censoring, where the probability of censoring is related to the event of interest, these metrics can give biased estimates of the underlying model error. In this paper, we introduce three new evaluation metrics for survival analysis based on Archimedean copulas that can account for dependent censoring. We also develop a framework to generate realistic, semi-synthetic datasets with dependent censoring to facilitate the evaluation of the metrics. Our experiments in synthetic and semi-synthetic data demonstrate that the proposed metrics can provide more accurate estimates of the model error than conventional metrics under dependent censoring.
1 Introduction
Learning survival prediction models resembles learning regression models, as both are essentially trying to learn, from a labeled dataset, a model that maps a description of an instance to a real value. Survival prediction examples include predicting the time to death in patients with amyotrophic lateral sclerosis (ALS) [1], predicting the time to a fall in older adults [2], or evaluating a new experimental treatment for cancer [3]. However, survival prediction differs from regression as some of the training instances are “censored” – that is, we only know a lower bound of the true time. This includes, for example, individuals who drop out or are still alive at the end of the study. The independent censoring assumption states that, within any subgroup of interest, censored patients at time $t$ are representative of all at-risk patients at that time with respect to their survival experience [4, Ch. 1]. Formally, we have $T\perp\mkern-9.5mu\perp C$, where $T$ is the event distribution and $C$ is the censoring distribution. This can also be stated explicitly in terms of the features $X$ as the conditional independent censoring assumption, i.e. $T\perp\mkern-9.5mu\perp C\mid X$ [5, Ch. 1] (see Appendix B for more details). Many standard survival tools – e.g., the Kaplan-Meier (KM) [6] and Nelson-Aalen [7, 8] estimators, and the Cox Proportional Hazards (CoxPH) model [9] – rely on the assumption of independent censoring [5, Ch. 1]. In the absence of features, if any confounder related to the event censors the event time, we refer to this phenomenon as dependent censoring.
Figure 1: (a) Residual dependence between event time $T$ (cancer relapse) and censoring time $C$ (cancer death) denoted by the diamond “Dep”, due to an unobserved confounding covariate – here “Tumor grade”. (b) Evaluation under dependent censoring, where Patient A is censored immediately before the event for reasons related to the event. This leads to a biased estimation of the marginal survival distribution using the Kaplan-Meier (KM) [6] estimator. Note that the two curves coincide between the range $[0,c_{i}]$. (c) Scatter plot generated from a standard exponential distribution under an Archimedean Clayton [10] copula.
Figure 1(a) shows an example of dependent censoring by omitting a confounding covariate – e.g.“tumor grade”. Since “tumor grade” relates to both $T$ (cancer relapse) and $C$ (cancer death), the variation in the covariate induces changes in both $T$ and $C$. That is, we want a model that predicts the time until relapse, but death is censoring. As “tumor grade” is predictive of tumor progression, an advanced (resp., early) stage is linked to shorter (longer) values for both cancer relapse $T$ and cancer death $C$, and consequently, $T$ and $C$ are not independent. This shows that the conditional independent censoring assumption might not hold if any confounding covariates are omitted or ignored in a regression model [4, Ch. 1].
Figure 1(b) demonstrates how dependent censoring can bias the marginal survival estimates from the KM estimator [6]. Consider patient A, censored (due to cancer death) exactly one year after study entry, who likely would have experienced cancer relapse shortly thereafter. However, the KM estimator may overestimate this patient’s survival time by assuming that they are like those remaining in the study, who tend to live longer, because it does not consider that these censored patients tend to have shorter event times [11]. To overcome this limitation, we need a survival estimator that can account for dependent censoring when predicting the marginal distribution. To this end, Zheng and Klein [12] generalized the KM estimator to the Copula-Graphic (CG) estimator under an Archimedean copula, $C_{\theta}$, which leverages assumptions about the dependence structure between event and censoring. Figure 1(c) illustrates sampling from such a copula. In this example, we see lower tail dependence from the copula – describing the probability that small (early) values of one variable (e.g., event) are associated with small values of another variable (e.g., censoring). The CG estimator then uses this information when estimating the population-based survival probability.
To motivate our approach, it is important to distinguish model estimation and model evaluation. Under the (conditional) independence assumption, we can reduce dependent censoring at the estimation stage by simply collecting more data (i.e., features) and obtaining model parameters using e.g., standard CoxPH [9, 13, 14], Random Survival Forest (RSF) [15] or Multi-Task Logistic Regression (MTLR) [16]. Under the independence assumption, we can use recent copula-based techniques to jointly learn the parameters of the marginal distributions and the copula $C_{\theta}$ between them [17, 18, 19]. However, these approaches do not account for dependent censoring at the estimation stage, since standard survival metrics – such as Harrell’s concordance index (CI) [20], the Brier Score (BS) [21], and the mean absolute error (MAE) using one of hinge [22], margin [22] or pseudo times [23] – still assume that $T$ and $C$ are independent, regardless of the model parameters. If $T$ and $C$ are indeed dependent, assuming independence can lead to inaccurate performance estimates [4, Ch. 1].
Contribution.
To overcome these limitations, we introduce three new metrics – CI-Dep, IBS-Dep, and MAE-Dep – to evaluate survival models in the presence of dependent censoring. These new metrics can evaluate the concordance (i.e., ranking) of the predicted risk scores, the accuracy of the survival curve, and the accuracy of the time-to-event estimates, respectively. They achieve this by accounting for the dependency between event and censoring times using an Archimedean copula. We also design a framework to create realistic, semi-synthetic datasets, with dependent censoring to evaluate the existing and proposed metrics. We show empirically that our metrics can provide more accurate estimates of the model error than existing metrics, however, the performance ultimately depends on the dependence structure in the data and the censoring rate. To our knowledge, these are the first survival metrics that explicitly account for dependent censoring.
Our code is publicly available at: https://github.com/thecml/DependentEVAL/
2 Background and related work
2.1 Survival data
Many survival analysis models estimate the probability that a specific event will occur for an individual at a time $T$ later than $t$, for all times $t>0$. Most often, the observation for a given instance will consist of a random variable, $T$, representing the time from a given origin ($t=0$) to the occurrence of the event, e.g.“death”. The distribution of $T$ can be characterized by the probability distribution function $F_{T}(t)=\text{Pr}(T\leq t)$ or, equivalently, by the survival function $S_{T}(t)=1-F_{T}(t)=\text{Pr}(T>t)$. Note that $S_{T}(t)$ (resp., $F_{T}(t)$) correspond to the probabilities of being event-free (resp., having experienced the event) by time $t$. We can estimate $S_{T}\left(t\right)$ nonparametrically, e.g., using the Kaplan-Meier [6] estimator (see Appendix G for a detailed derivation).
Definition 1.
The Kaplan-Meier (KM) estimator
$\begin{split}\hat{S}_{T}(t)\ &=\prod_{t_{i}\leq t,d_{i}=1}\left(1-\frac{1}{n_{i}}\right),\quad\hbox{\rm for}\quad t\leq t_{\text{max}}\end{split}$ (2.1)
where $n_{i}=\sum_{\ell=1}^{n}I(t_{\ell}\geq t_{i})$ is the number at-risk at time $t_{i}$; $\hat{S}(t)=1$ if no death occurs up to time $t$; $\hat{S}(t)$ is undefined for $t>t_{\text{max}}$ [6].
If features are entered into the model, we may define our survival dataset $\mathcal{D}$ as a set of observations, i.e., we have $\mathcal{D}=\left\{\left(\bm{x}_{i},t_{i},\delta_{i}\right)\right\}_{i=1}^{N}$, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a feature (covariate) vector for instance $i$, $t_{i}\in\mathbb{R}^{\geq 0}$ is the time of either censoring or event, depending on which occurred first, and $\delta_{i}\in\{0,1\}$ is a event indicator where $\delta_{i}=0$ means the event has not happened by $t_{i}$ (right-censored) and $\delta_{i}=1$ means the event occurred at $t_{i}$ (uncensored). We consider each instance to have a potential event time $e_{i}$ and a right-censoring time $c_{i}$, and observe only the earlier of the two, so the observed time and event indicator are defined as $t_{i}:=\min\{e_{i},c_{i}\}$ and $\delta_{i}:=\mathbbm{1}[e_{i}\leq c_{i}]$, where $\mathbbm{1}[\cdot]$ denotes the indicator function. Given survival data $\mathcal{D}$, let $f_{T}=\text{Pr}(T=t)$ denote the density, and $S_{T}=\text{Pr}(T>t)$ the survival function, then we have the likelihood under (conditional) independent censoring:
$\mathcal{L}(\mathcal{D})=\prod_{i=1}^{N}\left[f_{T}(t_{i}\mid\bm{x}_{i})\,S_{C}(t_{i}\mid\bm{x}_{i})\right]^{\delta_{i}}\left[f_{C}(t_{i}\mid\bm{x}_{i})\,S_{T}(t_{i}\mid\bm{x}_{i})\right]^{1-\delta_{i}}.$ (2.2)
2.2 Copula-based estimation
A copula is a function that links two or more random variables by specifying their dependence structure [4, Ch. 1]. Among the various types of copulas, the Archimedean copula family is the most used, and specifies the dependence structure as a single parameter, $\theta$. This family of copulas is defined by a generator function, $\varphi_{\theta}(t):[0,1]\to[0,\infty]$, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [4, Ch. 3]. Here, we will focus on two common Archimedean copulas: Clayton [10] and Frank [24]. Zheng and Klein [12] generalized the KM estimator (Equation 2.1) under independent censoring to the Copula-Graphic (CG) estimator under dependent censoring, which has become an important tool for analyzing survival data [25, 26, 27]. Rivest and Wells [28] obtained the first explicit expression of the CG estimator under an assumed copula from the Archimedean family (e.g., Clayton or Frank). We can estimate the survival function using the CG estimator (see Appendix G for a detailed derivation):
Definition 2.
The Copula-Graphic (CG) estimator
$\begin{split}\hat{S}_{T}(t)=\varphi^{-1}\left(\sum_{{i:\,}t_{i}\leq t,\delta_{i}=1}\varphi\left(\frac{n_{i}-1}{n}\right)-\varphi\left(\frac{n_{i}}{n}\right)\right),\quad\hbox{for}\quad t\ \leq\ t_{\text{max}}\end{split}$ (2.3)
where $n_{i}=\sum_{\ell=1}^{n}I(t_{\ell}\geq t_{i})$ is the number at risk at time $t_{i}$; $\hat{S}_{T}(t)=1$ if no death occurs up to time $t$; $\hat{S}_{T}(t)$ is undefined for $t>t_{\text{max}}$ [4, Ch. 4]. Note that when $\varphi(\cdot)$ is uniform, this reduces to the standard KM estimator.
If features are entered into the model and given some Archimedean copula (e.g., Clayton or Frank), we can rewrite Equation 2.2 as the log-likelihood under dependent censoring (see Appendix E.3):
$\mathcal{L}(\mathcal{D})=\sum_{i=1}^{N}\delta_{i}\;\text{log}\left[f_{T}(t_{i}|\bm{x}_{i})\right]\;+\;\delta_{i}\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T}(t_{i}\rvert\bm{x}_{i})}_{u_{2}=S_{C}(t_{i}\rvert\bm{x}_{i})}\right]$
$+\;(1-\delta_{i})\;\text{log}\left[f_{C}(t_{i}|\bm{x}_{i})\right]+(1-\delta_{i})\;\text{log}\left[\frac{\partial}{\partial u_{2}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T}(t_{i}\rvert\bm{x}_{i})}_{u_{2}=S_{C}(t_{i}\rvert\bm{x}_{i})}\right]$ (2.4)
In this expression, the first term represents the log-likelihood of observing the event at time $t_{i}$. The second term captures the conditional probability of the censoring time occurring after the event time, given that the event occurred at $t_{i}$. By symmetry, the third and fourth terms mirror these quantities but with respect to the censoring time. Although Equation 2.2 may appear complex, the partial derivatives of any Archimedean copula have closed-form expressions, which allows us to compute the log-likelihood function using gradient descent [17].
2.3 Evaluation metrics and limitations
We now outline the conventional evaluation metrics in survival analysis and discuss why these metrics are problematic under dependent censoring. See Appendix D for their formal definitions.
Concordance index (CI) evaluates the ranking of individuals when their relative ordering is of primary interest. It measures the proportion of concordant pairs – pairs where the predicted ranking of risk scores aligns with the true ordering of event times – among all comparable pairs [29]. A pair $(i,j)$ is defined as comparable if $\delta_{i}=1$ and $t_{i}<t_{j}$, which means that the earlier individual has experienced the event and the observed time of the later individual is greater than the first. Even under the random censoring assumption, it is evident that the comparable pairs selected using the above criteria do not represent all possible pairs based on their true event times:
$\text{Pr}(\delta_{i}=1,t_{i}<t_{j})\quad\neq\quad\text{Pr}(e_{i}<e_{j}).$
This inequality arises because the left-hand term can be decomposed as:
$\text{Pr}(\,\delta_{i}=1,\ t_{i}<t_{j})$ $=\quad\text{Pr}(\,e_{i}<c_{i},\ \ e_{i}\ <\ \min\{c_{j},e_{j}\}\ )$
$=\quad\text{Pr}(\,e_{i}<c_{i},\ t_{e,i}<c_{j},\ e_{i}<e_{j})$
$=\quad\underline{\text{Pr}(\,c_{i}<c_{i})\ \text{Pr}(e_{i}<c_{j})}\quad\text{Pr}(e_{i}<e_{j})\ ,$
where the factorization follows from the independent censoring assumption. Compared to the right-hand side term, this expression includes additional underlined terms related to the censoring distribution, confirming the findings of [30]. To address this limitation, Uno et al. [30] introduced a modified version of CI that employs inverse probability of censoring weights (IPCW), ensuring consistency with the true CI. The IPCW approach estimates the probability of being uncensored at each time point and reweighs the uncensored instances accordingly.
Integrated Brier Score (IBS) evaluates the overall accuracy of the predicted survival functions by integrating the Brier Score (BS) [21] over a range of time points. It measures the average squared difference between the predicted survival probabilities and actual survival outcomes over time [31]. To handle censored instances, IPCW can be used to reweigh the uncensored instances when computing both BS and IBS [32]. Similar to CI-IPCW [30], IPCW weights are often approximated using the KM estimator, which is biased under dependent censoring [4, Ch. 1].
Mean Absolute Error (MAE) quantifies the accuracy of point predictions for event times by measuring the absolute difference between the predicted and true event times. For censored instances, Haider et al. [22] proposed an approach that approximates that subject’s event time as the mean of the KM estimator (Equation 2.1), renormalized at that individual’s censoring time. This proxy event time, weighted by confidence estimates (also derived from KM), is then used to compute the absolute error. We will refer to this method as MAE-Margin. Given a subject is censored at time $t_{i}$, using some estimator $\hat{S}(t)$, e.g., the KM, the subject’s margin time is calculated as:
[…]
Limitations. All the aforementioned metrics typically rely on KM estimates to approximate either the marginal time-to-event or the marginal time-to-censoring distribution. While KM is advantageous due to its consistency under the independent censoring assumption, it becomes problematic when censoring and event times are dependent [4, Ch. 1]. In such scenarios, the KM estimates may be biased and inconsistent [33]. To mitigate these issues, researchers have proposed modifying these metrics by replacing KM with a conditional estimator (e.g.,  CoxPH) to account for the dependency between covariates and censoring distributions [31]. However, this approach still has two major limitations: (1) model misspecification: if the conditional estimator is inaccurately specified, the resulting estimates may be unreliable, and (2) uncaptured dependencies: under dependent censoring (i.e.,  when unobserved confounders exist), even a well-specified conditional estimator may fail to fully account for the dependency if there are important covariates missing in the analysis. These challenges highlight the need for robust estimation methods that can better handle dependent censoring scenarios.
3 The proposed metrics
CI-Dep: We propose a modified version of Harrell’s CI [20] that employs inverse probability weighting (i.e., IPCW), but the marginal survival probability for censoring at time $t$ is estimated using the CG estimator under an approximated Archimedean copula, parameterized by $\hat{\theta}$. When $\hat{\theta}\approx 0$, it yields the independence copula, which is equivalent to using the KM estimator as in [30]. Our proposed concordance index thus requires a known or approximated Archimedean copula that models the dependency between marginals $T$ and $C$, but does not require interaction with the covariates. Given an estimated risk function $\hat{\eta}(\cdot)$ and corresponding labels, the traditional CI is computed as:
$\begin{split}\text{CI}(\mathcal{D};\,\hat{\eta}(\cdot))=\frac{\sum_{i,j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]}.\end{split}$ (3.1)
Now, employing the IPCW technique used by [30], but under dependent censoring, let $\hat{G}_{\text{CG}}(t)=\text{Pr}(C>t)$ represent the marginal survival probability for censoring at $t$ estimated by the CG estimator. We then reweigh Equation 3.1 using the CG estimator as:
$\begin{split}\text{CI}_{\text{Dep}}(\mathcal{D};\,\hat{\eta}(\cdot))\ =\ \frac{\sum_{i,j}\delta_{i}\cdot\{\hat{G}_{\text{CG}}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j}\delta_{i}\cdot\{\hat{G}_{\text{CG}}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]},\end{split}$ (3.2)
which denotes the proposed CI-Dep. Here, the weight is given by $1/\hat{G}_{\text{CG}}(t)$.
IBS-Dep: We propose a modified version of the Brier Score [32] under dependent censoring using margin time estimates as the censoring weights. The margin or “best-guess” time can be interpreted as a conditional expectation of the event time given that the event time is greater than the censoring time. Given an instance is censored at time $t_{i}$, we can calculate its margin time using Equation 2.5, but with the CG estimator instead of the KM, i.e.$e^{\text{margin}}(t_{i})=e^{\text{margin}}(\hat{S}_{\text{CG}},t_{i})$, where $\hat{S}_{\text{CG}}$ is the marginal survival probability estimated using the CG estimator, then we have:
$\begin{array}[]{l}\text{BS}_{\text{Dep}}\left(\mathcal{D};\,\hat{S}(\cdot),t^{*}\right)=\frac{1}{N}\sum_{i=1}^{N}\left(\mathbbm{1}[\tilde{e}_{i}>t^{*}]-\hat{S}(t^{*}\mid\bm{x}_{i})\right)^{2},\end{array}$ (3.3)
$\text{where}\quad\tilde{e}_{i}=\begin{cases}e_{i}&\text{if}\quad\delta_{i}=1\\ e_{i}^{\text{margin (CG)}}(t_{i})&\text{elsewhere.}\end{cases}$
To calculate the expectation of Equation 3.3 over all times, we simply take the integral from $t=0$ to $t=t_{\text{max}}$
[…]
where $t_{\text{max}}$ is usually the maximum observed time in the training set.
MAE-Dep: We propose a modified version of the MAE-Margin [22] under dependent censoring. Again, we make use of the CG estimator to calculate margin times for the censored instances (Equation 2.5) and replace the censored event times with their “best-guess” estimates. Furthermore, for each censored instance, we adopt the approach from [22] and assign a confidence weight $\omega_{i}=1-\hat{S}_{\text{\text{CG}}}(t_{i})$ to the error based on the margin value estimated using the CG method. For uncensored subjects, we set $w_{i}=1$. The dependent MAE is thus:
$\text{MAE}_{\text{Dep}}\left(\mathcal{D};\,\hat{t}_{i}\right)=\frac{1}{\sum_{i=1}^{N}\omega_{i}}\sum_{i=1}^{N}\omega_{i}\Big{|}\big{[}(1-\delta_{i})\cdot e^{\text{margin (CG)}}(t_{i})+\delta_{i}\cdot t_{i}\big{]}-\hat{t}_{i}\Big{|}.$ (3.5)
4 Synthetic experiments
4.1 Experimental setup
We create synthetic datasets from a data generation process (DGP) under the Weibull distribution (see Appendix E.1). We sample R.V.s $T$ and $C$ from a copula given Kendall’s $\tau$; if $\tau=0$, the marginals are independent, otherwise if $\tau>0$, the marginals are from an Archimedean copula. In each of 100 experiments, we follow [34] and split the data into stratified training (70%), validation (10%), and test (20%) sets (seeds 0-99). Here, we investigate our metrics under dependency and censoring. See Appendix F.1 for evaluating under copula misspecification and varying sample sizes.
4.2 Results and takeaways
Dependent metrics under increasing dependency. Figure 2 shows the evaluation bias – i.e., the absolute difference between the ground-truth metric and the censored metric – for four baseline metrics (Harrell’s CI, Uno’s CI, IPCW-integrated Brier Score, and MAE-Margin) and our dependent variants as a function of Kendall’s $\tau$. Assuming a known ground-truth copula, we estimate $\hat{S}_{T}(t|\bm{x})$ using a CoxPH model to make survival predictions, where the risk scores are the negative predicted times. Overall, our dependent metrics show lower bias – i.e., closer proximity to the true error – when dependency is notably ($\tau>0.25$) than their independent variants. For the ranking metrics (Harrell’s CI, Uno’s CI), biases are similar under low dependency ($\tau<0.25$) but improve significantly under high dependency ($\tau>0.5$). For the accuracy metrics (IPCW-integrated Brier Score, MAE-Margin), our dependent variants give equal or slightly better error estimates regardless of the dependency.
Dependent metrics under increasing censoring. Figure 3 shows the evaluation bias as a function of the censoring rate (0-80%). Here, we set a fixed Kendall’s $\tau$ to 0.5, draw 10,000 samples from the DGP over 100 experiments, and group the results by the censoring rates in the whole dataset. We assume a known ground-truth copula and estimate $\hat{S}_{T}(t|\bm{x})$ using a CoxPH model. Our dependent metrics are generally aligned with their independent counterparts under low censoring (0-20%), but consistently exhibit lower biases and less variance when censoring increases ($\geq 20\%$).
Figure 2: Plot of mean biases ($\pm$ 2 SD.) for baseline and proposed metrics as a function of the dependency under a known copula, averaged over 100 experiments ($N=10,000$, $d=10$).
Figure 3: Plot of mean biases ($\pm$ 2 SD.) for baseline and proposed metrics as a function of the censoring rate under a known copula, averaged over 100 experiments (Kendall’s $\tau$ = 0.5, $N=10,000$, $d=10$) and two copula models (Clayton and Frank).
5 Semi-synthetic experiments
5.1 Experimental setup
To evaluate our metrics on real datasets, we require knowledge of the true event times, which is not feasible due to censoring. To address this, we construct realistic semi-synthetic datasets where the covariates, event times, and censoring distribution reflect those of real-world datasets. Inspired by [23], we estimate a censoring distribution $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$ from a real dataset $\mathcal{D}$, remove censored instances to obtain $\mathcal{D^{\prime}}$, and then synthetically censor instances in $\mathcal{D^{\prime}}$ based on whether a sampled censoring time is earlier than the event time ($\tilde{c}_{i}<e_{i}$). We induce dependent censoring by omitting features in the original dataset based on one of three feature selection strategies. This ultimately gives us $\mathcal{D}^{\prime\prime}$, which has a synthetic but realistic censoring distribution, a restricted feature set, and known ground-truth event times, enabling us to compare the proposed and existing (censored) metrics to the true metrics. Figure 4 illustrates the evaluation pipeline. The remainder of this section outlines the practical details of our approach.
Figure 4: Evaluation pipeline. (1) Based on some raw, survival dataset $\mathcal{D}$, (2) we form $\tilde{\mathcal{D}}$ by flipping the event bit of $\mathcal{D}$ and then estimate the feature-dependent censoring distribution, $G_{\text{CoxPH}(\tilde{\mathcal{D}})}$. (3) We then select a subset of features from $\mathcal{D}$ using some strategy (e.g., Top-K features), (4) and generate a ground-truth dataset, $\mathcal{D}^{\prime}$, by removing the censored instances from $\mathcal{D}$. (5) Putting all this together, we have a semi-synthetic dataset $\mathcal{D}^{\prime\prime}$, which is based on the ground-truth dataset, the learned censoring distribution, and the selected features. (6) We then train two Archimedean copula models to estimate $\hat{C}_{\theta}$ and (7) five survival learners to make predictions based on $\mathcal{D}^{\prime\prime}$. Finally, (8) we compare the baseline and proposed censored metrics (using $\hat{C}_{\theta}$) with the true evaluation metric.
Datasets (Step 1 in Figure 4): We apply the evaluation pipeline to 6 large, real-world datasets: METABRIC [35], 2 from MIMIC-IV (all causes and hospital death) [36], and 3 from SEER (death from brain, liver and stomach cancer) [37]. They differ in their number of instances ($N$), number of features ($d$) and censoring rates ($C$). See Appendix E.1 for more details.
Feature selection (Step 3 in Figure 4): To induce dependent censoring, we apply one of three feature selection strategies to $\mathcal{D}$: (1) Original, where all features are retained; (2) Top-$k$, where we select the $k$ most predictive features based on permutation feature importance [38] evaluated using Harrell’s CI [20]; and (3) Random-%, where a random subset of features is selected.
Copula model (Step 6 in Figure 4): We estimate $\hat{C}_{\theta}$ by training Clayton and Frank copulas with nonlinear Weibull proportional-hazards marginals for $T$ and $C$ on $\mathcal{D}^{\prime\prime}$. Each marginal uses a multilayer perceptron (MLP) with one hidden layer and ReLU activation function. Marginals and copulas are trained jointly via MLE (see Appendix E.3). The copula with the lowest validation NLL is selected for evaluation.
Survival learners (Step 7 in Figure 4): We compare the prediction results using 5 survival learners produced by the CoxPH [9], Gradient Boosting Survival Analysis (GBSA) [39], Random Survival Forests (RSF) [15], DeepSurv [13] and Multi-Task Logistic Regression (MTLR) [16]. These learners are trained to predict an individual survival distribution (ISD) for each instance in the test set. See Appendix E.2 for further details, including implementation details and hyperparameters.
We repeat the procedure outlined in Figure 4 over 5 experiments. In each of 5 experiments, we follow [34] and split $\mathcal{D}^{\prime\prime}$ into stratified training (70%), validation (10%), and test (20%) sets (seeds 0-4). After splitting, we impute missing values using the sample mean for real-valued covariates or the mode for categorical covariates based on the training set only. We then normalize numerical values and encode categorical features using a one-hot encoding strategy.
5.2 Results and takeaways
Due to space constraints, the main text presents only partial results for datasets with different characteristics, e.g., number of samples and censoring rates. See Appendix F.2 for the complete results, where we report the approximated copula $\hat{C_{\theta}}$ and also evaluate the ranking accuracy.
Discrimination. Table 1 presents the discrimiation (CI) results in terms of the evaluation bias in 3 semi-synthetic datasets: METABRIC, MIMIC-IV (all), and SEER (liver). In the METABRIC dataset, CI-Dep performs very similarly to Uno’s CI on Original, Top 5, and Top 10 strategies (bias $\approx$ 0.005) and slightly higher than Harrell’s CI. In the MIMIC-IV dataset, all the metrics are generally aligned with similar mean biases and error margins. In the SEER dataset, CI-Dep slightly diverges from CI-Uno, giving marginally higher bias on Top 10 and lower on Rand. 25% strategies. Overall, we see that CI-Dep does not significantly improve or degrade performances over the baseline metrics.
Accuracy. Table 2 presents the accuracy (IBS and MAE) results in terms of the evaluation bias the same datasets. In the METABRIC dataset, all IBS results are identical or near-identical across metrics and strategies and the MAE-Dep has slightly lower bias than Margin, which is its independent counterpart. In the MIMIC-IV dataset, IBS-Dep performs on par with KM, and MAE-Dep again closely matches other metrics, except Hinge. In the SEER dataset, IBS-Dep equals KM and IPCW, but the MAE-Dep performs better than Hinge, Pseudo and Margin in Top 5 and Top 10 strategies. We hypothesize that the similar biases are likely due to a lack of dependent censoring or the relatively low censoring rates (0–20%). In our synthetic experiments, more pronounced differences between the independent and dependent metrics typically emerged beyond 25% censoring (see Figure 3).
Table 1: Comparison of baseline and proposed discrimination metrics (CI) in terms of the mean ($\pm$ 2 SD.) bias, i.e., the proximity to the true error, averaged over 5 survival learners and 5 experiments.
Dataset Strategy Harrell Uno Dep.
METABRIC ($N$=1,102, $C$=3.8%) Original 0.001$\pm$0.002 0.004$\pm$0.003 0.005$\pm$0.005
Top 5 0.001$\pm$0.002 0.005$\pm$0.003 0.006$\pm$0.005
Top 10 0.001$\pm$0.002 0.004$\pm$0.002 0.005$\pm$0.003
Rand. 25% 0.001$\pm$0.002 0.023$\pm$0.014 0.023$\pm$0.014
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original 0.006$\pm$0.004 0.007$\pm$0.004 0.007$\pm$0.004
Top 5 0.011$\pm$0.004 0.016$\pm$0.005 0.016$\pm$0.005
Top 10 0.010$\pm$0.003 0.013$\pm$0.004 0.013$\pm$0.004
Rand. 25% 0.006$\pm$0.002 0.006$\pm$0.003 0.006$\pm$0.003
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.008$\pm$0.002 0.006$\pm$0.003 0.008$\pm$0.003
Top 5 0.019$\pm$0.002 0.017$\pm$0.003 0.017$\pm$0.003
Top 10 0.008$\pm$0.002 0.009$\pm$0.002 0.011$\pm$0.001
Rand. 25% 0.011$\pm$0.014 0.019$\pm$0.017 0.018$\pm$0.017
Table 2: Comparison of baseline and proposed accuracy metrics (IBS and MAE) in terms of mean ($\pm$ 2 SD.) bias, i.e., the proximity to the true error, averaged over 5 survival learners and 5 experiments. KM denotes the proposed IBS metric under the independent copula for comparison. Bold indicates best performing independent/dependent variant for the MAE.
IBS MAE
Dataset Strategy IPCW KM Dep. Hinge Pseudo Margin Dep.
METABRIC ($N$=1,102, $C$=3.8%) Original 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.246$\pm$0.414 0.314$\pm$0.459 0.328$\pm$0.431
Top 5 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.249$\pm$0.398 0.308$\pm$0.460 0.298$\pm$0.464
Top 10 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.248$\pm$0.418 0.317$\pm$0.462 0.308$\pm$0.465
Rand. 25% 0.012$\pm$0.018 0.012$\pm$0.018 0.012$\pm$0.018 1.397$\pm$0.416 0.266$\pm$0.435 0.342$\pm$0.463 0.328$\pm$0.467
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original 0.008$\pm$0.003 0.009$\pm$0.003 0.009$\pm$0.003 62.244$\pm$10.539 37.156$\pm$10.255 37.156$\pm$10.255 37.301$\pm$10.245
Top 5 0.008$\pm$0.004 0.011$\pm$0.004 0.011$\pm$0.004 65.626$\pm$11.943 39.558$\pm$11.741 39.558$\pm$11.741 39.717$\pm$11.743
Top 10 0.008$\pm$0.004 0.010$\pm$0.004 0.010$\pm$0.004 63.716$\pm$11.380 38.131$\pm$11.190 38.131$\pm$11.190 38.312$\pm$11.205
Rand. 25% 0.008$\pm$0.003 0.011$\pm$0.004 0.011$\pm$0.004 62.749$\pm$8.117 37.204$\pm$8.102 37.204$\pm$8.102 37.364$\pm$8.062
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.003$\pm$0.003 0.003$\pm$0.002 0.003$\pm$0.002 2.498$\pm$0.075 2.475$\pm$0.134 1.478$\pm$0.110 1.520$\pm$0.107
Top 5 0.003$\pm$0.003 0.003$\pm$0.002 0.003$\pm$0.002 2.688$\pm$0.122 2.246$\pm$0.134 1.250$\pm$0.123 1.226$\pm$0.219
Top 10 0.003$\pm$0.003 0.002$\pm$0.002 0.002$\pm$0.002 2.513$\pm$0.072 2.462$\pm$0.119 1.466$\pm$0.102 1.450$\pm$0.170
Rand. 25% 0.005$\pm$0.005 0.004$\pm$0.004 0.004$\pm$0.004 2.505$\pm$0.136 2.499$\pm$0.272 1.490$\pm$0.193 1.524$\pm$0.227
6 Discussion
Limitations. This work has limitations: First, our metrics rely on a known or well-approximated Archimedean copula (see the following paragraph) to adjust for dependent censoring. However, it is typically impossible to recover the true underlying copula due to the nature of competing-risks data: for each instance, we only observe one outcome (either event or censoring), but not both [40]. This is a key limitation of applying our metrics and of copula-based methods in general, since the ground-truth dependence structure cannot be directly observed or learned. Second, our metrics may be misspecified under a temporal distribution shift in the dependence structure. That is, if $P(T\mid X)$, $P(C\mid X)$, and $P(T,C\mid X)$ change over time, then the dependence structure evolves, and we would have to re-estimate the copula parameter. Third, successfully maximizing the likelihood (Equation 2.2) depends on the dataset characteristics – e.g., number of instances, censoring rate, etc. – which means that model estimates may have high variance in small sample sizes.
User guide. Our metrics require users to specify a dependence structure under an Archimedean copula. However, since no nonparametric test can conclusively detect the dependency between event and censoring, we instead offer a practical approach by estimating a likely dependence structure based on the log-likelihood: the question is not whether this will yield the optimal (latent) structure, but a structure better than simply assuming independence between event and censoring. The log-likelihood is a widely used loss function in survival analysis and is also a proper scoring rule [41]. We first establish a baseline using the independent copula model, then assess whether a dependent copula provides a better fit by approximating the data’s dependence structure. As a useful heuristic, $\mathcal{D}$ should contain 25% censoring or more, based on our observations, for dependent censoring to be relevant.
  1. 1.
    Given $\mathcal{D}$, split the data into stratified training, validation, and test sets. Define marginals (e.g., Weibulls) for event and censoring and maximize the dependent-likelihood (Equation 2.2) under a Clayton independent-copula $C_{\theta}$ based on the training data. This can be done using gradient descent, setting $\theta$ to a small positive constant (e.g.$10^{-4}$ such that $\theta>0$) and freezing its weight in the optimization procedure (see Appendix E.3). Record the validation loss and halt training when it stops improving. This will yield survival marginals, a fixed copula $\hat{C_{\theta}}\approx 0$ under independence, and some validation loss $\mathcal{L}_{\text{Indep}}$, which is our baseline.
  2. 2.
    Reset the marginal model weights and reinitialize $C_{\theta}$ as an Archimedean copula, but unfreeze its weight in the optimization procedure and maximize the dependent-likelihood (Equation 2.2) based on the training data. Again, record the validation loss and halt training when it stops improving. This will again yield survival marginals, a copula $\hat{C_{\theta}}$ under dependence, and some validation loss $\mathcal{L}_{\text{Dep}}$. Repeat Step 2 for each copula model to be evaluated.
  3. 3.
    Now, compare $\mathcal{L}_{\text{Indep}}$ and $\mathcal{L}_{\text{Dep}}^{k}$ for each of the $K$ trained copula models under dependence. If $\mathcal{L}_{\text{Dep}}^{k}<\mathcal{L}_{\text{Indep}}$, then the $k$-dependent model better captures the joint distribution, implying a better model fit. Alternatively, use a penalized likelihood criterion (e.g., AIC, BIC) and choose the model with the lower score to balance goodness-of-fit and complexity. Finally, use the approximated $\hat{C_{\theta}}$ from the chosen model to calculate our metrics. Recall that our metrics reduce to their independent variants when $C_{\theta}\approx 0$.
Broader impacts. Benchmarking survival models is inherently difficult due to censoring, as we do not have complete information about the survival times for some instances. Making wrong assumptions about the censoring mechanism can give a false impression of how well a model actually performs. This work advances the reliable evaluation of survival models under dependent censoring, which is important for real-world applications – such as biomedical and epidemiological research – where accurate predictions inform clinical decisions and public health policy.
7 Conclusion
In this paper, we have proposed three copula-based metrics for evaluating survival models under dependent censoring. We empirically compared them to existing, independent variants in both synthetic and semi-synthetic datasets. We found through empirical evaluation that our metrics can more accurately estimate model error than existing alternatives, though their effectiveness depends on the dependence structure in the data and the amount of censoring. We anticipate that others will be able to use these metrics and improve upon our framework.
Acknowledgments and Disclosure of Funding
This research received support from the Natural Science and Engineering Research Council of Canada (NSERC), the Canadian Institute for Advanced Research (CIFAR), and the Alberta Machine Intelligence Institute (Amii).
Appendix A Notation
Table 3 provides a list of notations used in this work.
Table 3: Table of notation.
Symbol Definition
$c_{i}$ Censoring time of instance $i$
$e_{i}$ Event time of instance $i$
$t_{i}$ Observed time of instance $i$
$\hat{t}_{i}$ Predicted time of instance $i$
$\delta_{i}$ Event bit/indicator, $\delta_{i}=\mathbbm{1}[e_{i}>c_{i}]$
$f_{T}$ Probability density function, representing $\text{Pr}(T=t)$
$F_{T}$ Cumulative density function, representing $\text{Pr}(T<t)$
$S_{T}$ Survival function, $S:\mathbb{R}\rightarrow[0,1]$
$\mathcal{D}$ Raw dataset
$\mathcal{D^{\prime}}$ Raw dataset with only uncensored subjects
$\mathcal{D^{\prime\prime}}$ Semi-synthetic dataset with synthetic censoring on $\mathcal{D^{\prime}}$
$\tilde{\mathcal{D}}$ Raw dataset with flipped event bit
$N$ Number of observations in the dataset
$d$ Number of features/covariates in the dataset
$C$ Censoring rate (%) in the dataset
$\bm{x}_{i}$ Covariates of instance $i$
$G(t)$ Censoring distribution
$G_{CoxPH}(t\mid\bm{x})$ Feature-dependent censoring distribution, estimated using CoxPH
$h_{0}(t)$ Baseline hazard function
$h(\cdot\mid\bm{x})$ Hazard function given the covariates $\bm{x}$
$\mathcal{L}$ Objective/loss function
$C_{\theta}$ Copula parameterized by $\theta$
$u_{1},u_{2}$ Inputs to a copula function
Appendix B Censoring assumptions
Random censoring. We define random censoring as follows: “Subjects who are censored at time t should be representative of all the study subjects who remained at risk at time t with respect to their survival experience.” [5, Ch. 1]. In other words, we assume that the death rate (or failure rate) of censored subjects is the same as that of uncensored subjects who remain in the risk set. Random censoring is a stronger assumption and more restrictive than what is called independent censoring, which states that censoring is independent within a subgroup.
Definition 3.
Random censoring
$h_{\text{CE}}(t)=h_{\text{NCE}}(t),$ (B.1)
where $h_{\text{CE}}$ is the hazard rate for the censored subjects, and $h_{\text{NCE}}$ is the hazard rate for the uncensored subjects.
Independent censoring. In the absence of competing risks, we define independent censoring as follows: “Within any subgroup of interest, subjects who are censored at time $t$ should be representative of all subjects in that subgroup who remained at risk at time $t$ with respect to their survival experience” [5, Ch. 1]. In other words, censoring is independent provided that it is random within any subgroup given the covariates.
Definition 4.
Independent censoring
$h_{A,\text{CE}}(t)$ $=h_{A,\text{NCE}}(t),$ (B.2)
$h_{B,\text{CE}}(t)$ $=h_{B,\text{NCE}}(t),$
where $A$ and $B$ are subgroups, e.g., patients given treatment A and patients given treatment B.
The idea behind random and independent censoring is that, even though the censored subjects were not randomly selected, their survival experience would be expected to be the same as if they had been randomly selected from the risk set at time $t$. Consider a clinical trial studying the recovery time of patients after surgery. The patients’ recovery times are measured, but some individuals leave the study early for reasons unrelated to their recovery, such as relocating to another city, personal commitments, or withdrawal for reasons not related to their health (e.g., family obligations). These individuals are censored, but their reason for leaving the study has nothing to do with the outcome (recovery time). Many of the analytic techniques in survival analysis, e.g., the Kaplan-Meier [6] and Nelson-Aalen [7, 8] population estimators, and the semiparametric Cox Proportional Hazards (CoxPH) model [9], rely on the assumption of independent censoring for valid inference in the presence of right-censored data [5, Ch. 1]. When censoring is not independent, however, this can introduce bias in the analysis. Such bias can arise when censored individuals have a higher likelihood of experiencing the event compared to those who are not censored. Consequently, if a significant portion of censored individuals had the event (e.g., died), the estimated survival probability at any time point may overestimate the true survival probability at that time. By example of [5, Ch. 1], consider a drug study where some patients are censored due to the occurrence of side effects. The unobserved survival outcomes of those censored because of side effects may not be representative of those who remain in the study. If patients who experience side effects are more susceptible to the health outcome, e.g., death, assuming independent censoring could lead to an overestimation of their survival rate.
Non-informative censoring. Censoring can be non-informative if the distribution of survival times ($T$) provides no information about the distribution of censoring times ($C$), and vice versa. Otherwise, the censoring is informative. However, we would still need to know which subjects are censored or not. The assumption of non-informative censoring is often justifiable when censoring is independent and/or random [5, Ch. 1], but cases exist where the censoring is informative but also random and independent: Consider a cohort study, where each time a patient experiences an event (e.g., death), another patient in the study is randomly selected to be censored. This could be a family member who decides to leave the study after an event occurs. If those who are censored are representative of those who remain in the risk set, then the censoring would be random and independent. However, in this case, the censoring would also be informative, as the censoring mechanism would be linked to the time-to-event distribution (since events trigger censoring). In fact, if this were the only mechanism by which individuals were censored, the distribution of survival times would entirely determine the distribution of censoring times, making it highly informative.
Appendix C Copulas and definitions
We can view bivariate survival analysis as a competing risks problem, where the occurrence of an event (e.g., death) and censoring (e.g., dropout) are mutually-exclusive incidents. Tsiatis [40] referred to this as the identifiability issue; competing risks data do not allow one to observe two event times simultaneously (one event censors the other), and hence, the data may not identify the dependence structure between event times. Zheng and Klein [12] gave a partial solution to the identifiability problem of dependent censoring by an assumed copula between two event times. A copula is a function that links two or more random variables by specifying their dependence structure [4, Ch. 1]. A mathematician, Abe Sklar, first used the word “copula” in his study of probabilistic metric space [42]. In his paper, he gave a mathematical definition of copulas and established the most fundamental theorem about copulas, known as Sklar’s theorem [42]. It demonstrated that any joint cumulative density can be written in terms of a copula over the quantiles of its marginal cumulative densities. Among the various types of copulas, the Archimedean copula is the most used, and enables the modeling of dependencies in arbitrarily high dimensions using a single parameter. This family of copulas is defined by a convex generator function, $\varphi:[0,1]\to[0,\infty]$, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [4, Ch. 3]. Given event time $T$ and censoring time $C$, we can define an Archimedean copula as:
$\begin{split}P(T>t,C>c)=\varphi^{-1}\big{[}\varphi\{S_{T}(t)\}+\varphi\{S_{C}(c)\}\big{]},\end{split}$ (C.1)
where $\varphi$ is the generator function, and $S_{T}(t)=P(T>t)$ and $S_{C}(t)=P(C>c)$ are the marginal survival functions. In the bivariate case, $\varphi$ describes the dependence structure (i.e., correlation) between two random variables, e.g.$T$ and $C$, and the (rank) correlation can be described with Kendall’s tau ($\tau$):
$\begin{split}\tau=&\quad\text{Pr}\{\ (t_{2}-t_{1})(c_{2}-c_{1})\geq 0\}-\text{Pr}\{\ (t_{2}-t_{1})(c_{2}-c_{1})<0\},\end{split}$ (C.2)
where ($t_{1}$, $c_{1}$) and ($t_{2}$, $c_{2}$) are sampled from a (survival) copula. The Clayton [10] and Frank [24] copulas are two common Archimedean copulas in survival literature. Within these families, the copula $C_{\theta}$ is parameterized by $\theta$, interpreted as the degree of dependence between ${T}$ and ${C}$ in the bivariate case. When $\theta\approx 0$, it yields the independence copula; thus, a larger value of $\theta$ implies greater dependency between the marginal distributions. Formally, we write $C(u_{1},\ldots,u_{d}):[0,1]^{d}\rightarrow[0,1]$, which is a $d$-dimensional copula (i.e., where $d$ is the number of variables) with $u_{1},\ldots,u_{d}$ uniform marginal probabilities and the following properties [43]:
1. Groundedness: If there exists an $i\in\{1,\dots,m\}$ such that $u_{i}=0$, then
$C(u_{1},\dots,u_{m})=0.$
2. Uniform Margins: For all $i\in\{1,\dots,m\}$, if $\forall j\neq i\Rightarrow u_{j}=1$, then
$C(u_{1},\dots,u_{m})=u_{i}.$
3. $d$-Increasingness: For all $u=(u_{1},\dots,u_{m})$ and $v=(v_{1},\dots,v_{m})$ where $u_{i}<v_{i}$ for all $i=1,\dots,m$, the following holds:
$\sum_{\ell\in\{0,1\}^{m}}(-1)^{\ell_{1}+\dots+\ell_{m}}C\left(u_{1}^{\ell_{1}}v_{1}^{1-\ell_{1}},\dots,u_{m}^{\ell_{m}}v_{m}^{1-\ell_{m}}\right)\geq 0.$
Let $T$ be the event distribution, $C$ the censoring distribution, $\bm{x}$ a vector of covariates, and $S_{T\mid\bm{x}}(t\mid\bm{x})=\text{Pr}(T>t\mid\bm{x})$ and $S_{C\mid\bm{x}}(t\mid\bm{x})=\text{Pr}(C>c\mid\bm{x})$ the marginal survival functions given $\bm{x}$. We can then define a bivariate (survival) copula $C_{\theta}$ [43] that describes the degree of dependence between $T$ and $C$ as follows:
$\text{Pr}(T>t,C>c\mid\bm{x})\ =\ C_{\theta}\{\,S_{T}(t\mid\bm{x}),\,S_{C}(c\mid\bm{x})\,\}.$ (C.3)
Kendall’s tau ($\tau$) can be defined as:
$\begin{split}\tau=&\quad\text{Pr}\{\ (t_{2}-t_{1})(c_{2}-c_{1})\geq 0\mid\bm{x}\ \}-\text{Pr}\{\ (t_{2}-t_{1})(c_{2}-c_{1})<0\mid\bm{x}\ \},\end{split}$ (C.4)
where ($t_{1}$, $c_{1}$) and ($t_{2}$, $c_{2}$) are sampled from the model (C.3). Kendall’s $\tau$ can be solely expressed as a function of the copula $C_{\theta}$ by [4, Ch. 3]:
$\tau_{\theta}=4\int_{0}^{1}\int_{0}^{1}C_{\theta}(u,v)\,C_{\theta}(du,dv),$ (C.5)
which implies that Kendall’s $\tau$ does not depend on the marginal survival functions. An Archimedean copula is a type of copula that is defined using a generator function, which allows for a flexible way to model dependence structures between random variables, e.g.$T$ and $C$. Such a copula is defined as [4, Ch. 3]:
$C_{\theta}(u,v)\ =\ \varphi_{\theta}^{-1}\left(\,\varphi_{\theta}(u)+\varphi_{\theta}(v)\,\right),$ (C.6)
where the function $\varphi_{\theta}:[0,1]\rightarrow[0,\infty]$ is called a generator of the copula, which is continuous and strictly decreasing from $\varphi_{\theta}(0)\geq 0$ to $\varphi_{\theta}(1)=0$. The following Archimedean copulas are considered in this work:
  • The Independence copula: $C(u,v)=uv$.
  • The Clayton copula [10]:
    $C_{\theta}(u,v)=\left(u^{-\theta}+v^{-\theta}-1\right)^{-1/\theta},\quad\theta\geq 0.$
  • The Frank copula [24]:
    $C_{\theta}(u,v)=-\frac{1}{\theta}\log\left[1+\frac{(\exp(-\theta u)-1)(\exp(-\theta v)-1)}{e^{-\theta}-1}\right].$
    $\quad\theta\neq 0.$
Appendix D Existing evaluation metrics
D.1 Concordance Index
Harrell’s concordance index (CI) quantifies the proportion of concordant pairs among all comparable pairs [29]. Given an estimated risk function $\hat{\eta}(\cdot)$ and corresponding labels, CI is computed as:
$\text{CI}(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{\eta}(\cdot))\ =\ \frac{\sum_{i,j:i\neq j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j:i\neq j}\delta_{i}\cdot\mathbbm{1}[t_{i}<t_{j}]}.$ (D.1)
Some definitions of CI [44] incorporate tie-breaking rules for cases where risk scores are identical. Specifically, a score of 0.5 may be assigned to tied pairs in the numerator. However, for simplicity, we omit ties in our discussion.
As shown in [30], this calculation converges to a censoring-dependent probability:
$P(\ \hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})\mid e_{i}<\min(e_{j},c_{j})\ ),$
which differs from the intended probability:
$P(\ \hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})\mid e_{i}<e_{j}\ ),$
To address this bias, Uno et al. [30] proposed the IPCW-weighted CI, computed as [^1] The proposed IPCW-CI also comes with a truncation time $\tau$ as a parameter – which means we only focus on the accuracy of ranking within $[0,\tau]$. Here, for consistency with other CIs, we do not consider such a truncation time.:
$\text{CI}_{\text{IPCW}}(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{\eta}(\cdot))\ =\ \frac{\sum_{i,j:i\neq j}\delta_{i}\cdot\{\hat{G}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]\cdot\mathbbm{1}[\hat{\eta}(\bm{x}_{i})>\hat{\eta}(\bm{x}_{j})]}{\sum_{i,j:i\neq j}\delta_{i}\cdot\{\hat{G}(t_{i})\}^{-2}\cdot\mathbbm{1}[t_{i}<t_{j}]},$ (D.2)
where $\hat{G}(t)$ represents the marginal survival probability for censoring at time $t$, typically estimated using the Kaplan-Meier (KM) estimator. The inverse probability censoring weight (IPCW) is given by $1/\hat{G}(t)$. In cases where censoring probability depends on features (i.e., $\hat{G}(t\mid\bm{x}_{i})$), a survival model such as CoxPH can be used for estimation. These two metrics are implemented in the SurvivalEVAL package [45].
D.2 Integrated Brier Score
The Brier Score (BS) quantifies the accuracy of probabilistic survival predictions at a specific time $t^{*}$ [32]. It measures the squared error between the predicted survival probability and the observed survival status at $t^{*}$:
To incorporate the censored subjects in the calculation, BS also requires estimating the IPCW to uniformly transfer a censored subject’s weight (originally 1) to subjects with known status at that time [46]. Formally:
BS ($\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{S}_{T}(\cdot\mid\cdot),\hat{G}(\cdot),t^{*}\right)\ =$
$\ \frac{1}{N}\sum_{i}\left(\frac{\delta_{i}\cdot\mathbbm{1}[t_{i}\leq t^{*}]\cdot\hat{S}_{T}(t^{*}\mid\bm{x}_{i})^{2}}{\hat{G}(t_{i})}+\frac{\mathbbm{1}[t_{i}>t^{*}]\cdot(1-\hat{S}_{T}(t^{*}\mid\bm{x}_{i}))^{2}}{G(t^{*})}\right).$ (D.3)
The integrated Brier score (IBS) is the expectation of BS over all time, defined as:
$\text{IBS}\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{S}_{T}(\cdot\mid\cdot),G(\cdot)\right)\ =\ \frac{1}{t_{\text{max}}}\cdot\int_{0}^{t_{\text{max}}}\text{BS}\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{S}_{T}(\cdot\mid\cdot),G(\cdot),\tau\right)d\tau,$ (D.4)
where $t_{\text{max}}$ is usually the maximum event time of the combined training and validation datasets. This metric is implemented in the SurvivalEVAL package [45].
D.3 Mean Absolute Error
Survival prediction shares similarities with regression, as it estimates a continuous event time given a feature set $\bm{x}_{i}$. Therefore, mean absolute error (MAE) serves as a useful evaluation metric.
MAE-Margin [22] estimates a “best guess”-event time for censored subjects using the KM estimator [6]. For a subject censored at time $t_{i}$, the margin event time is computed as:
[…]
where $S_{\text{KM}}(t)$ is the KM estimation that is typically derived from the training dataset.
Additionally, for each censored subject, Haider et al. [22] suggested we use a confidence weight $\omega_{i}=1-S_{\text{KM}}(t_{i})$ for the error calculated based on the margin value. This weight $\omega_{i}$ yields lower confidence for early censoring subjects and higher confidence for late censoring data. Of course, we set the weights for uncensored subjects $i$ to $w_{i}=1$ as we have full confidence in those error calculations. The overall MAE margin after a re-weighting scheme is:
$\text{MAE}_{\text{margin}}\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{t}_{i}\right)=\frac{1}{\sum_{i=1}^{N}\omega_{i}}\sum_{i=1}^{N}\omega_{i}\left|[(1-\delta_{i})\cdot e^{\text{margin}}(t_{i})+\delta_{i}\cdot t_{i}]-\hat{t}_{i}\right|.$
This metric is implemented in the SurvivalEVAL package [45].
Appendix E Experimental details
E.1 Data generation and preprocessing
Table 4: Overview of the raw datasets. The datasets are sorted by name. The number of features before performing one-hot encoding, with brackets (the number of features after one-hot encoding).
Dataset #Instances #Features %Censored #Event Max Event $t$
METABRIC 1,902 9 (9) 42.1% 1102 355
[…]
MIMIC-IV (all causes) 38,520 91 (103) 66.7% 12845 4404
[…]
SEER (liver) 82,841 14 (14) 37.6% 51704 225
SEER (stomach) 100,360 14 (14) 43.4% 56807 226
In this section, we describe how the synthetic dataset is generated and how we preprocessed the raw survival datasets. Table 4 shows an overview of the raw datasets.
Synthetic: This paper utilizes a linear synthetic dataset with two competing events (i.e., event and censoring), both based on the parametric Weibull distribution, as discussed in [17]. The datasets are characterized by the conditional hazard function and conditional survival probability of an event (or censoring) as follows:
$h_{T}(t\mid\bm{x}_{i})$ $=\left(\frac{v}{\rho}\right)\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(\bm{x}_{i})\right),$ (E.1)
$S_{T}(t\mid\bm{x}_{i})$ $=\exp\left(-\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(\bm{x}_{i})\right)\right),$ (E.2)
where $v$ and $\rho$ represent the shape and scale parameters of the Weibull distribution, respectively. We use $v_{T}=4$, $v_{C}=6$, $\rho_{T}=17$, $\rho_{C}=19$. $g_{\Psi}(\bm{x})$ represents the risk function, which is a linear function in this work. The data generation process begins by initializing a copula with a predefined parameter $\theta$. We then sample $N$ sets of CDF observations from the copula, each containing two CDF probabilities (corresponding to each competing event, i.e., event or censoring). From these probabilities, we then calculate the corresponding times using the inverse function of (E.2). The data generation process is further described in Algorithm 2 in [17].
METABRIC: The Molecular Taxonomy of Breast Cancer International Consortium (METABRIC) dataset [35] contains survival information for breast cancer patients. This dataset includes a diverse range of feature sets that encompass clinical traits, expression profiles, copy number variation (CNV) profiles, and single-nucleotide polymorphism (SNP) genotypes. All these features are derived from breast tumor samples collected during the METABRIC trial. The dataset can be downloaded from (https://www.cbioportal.org/study/summary?id=brca_metabric), and it does not have any missing values.
MIMIC-IV: The Medical Information Mart for Intensive Care (MIMIC)-IV [36] contains critical care data from patients admitted to hospitals and intensive care units (ICU). We create a version using the MIMIC-IV database, which contains patients who are alive at least 24 hours after being admitted to ICU. Their date of death is derived from hospital records or state records, which means the cause of mortality is not limited to the reason for ICU admission. We follow the instructions in [23] to process the dataset, and the code is available in their GitHub repository: https://github.com/shi-ang/CensoredMAE.
SEER: The SEER Program dataset [37] is a comprehensive collection of cancer patient data from approximately 49% of the U.S. population. Our study focuses on three subsets: SEER-brain, SEER-liver, and SEER-stomach, containing data on patients diagnosed with brain, liver, and stomach cancers, respectively. The goal is to model the time from diagnosis to failure events, such as death or disease progression. We preprocess the dataset as in [47]. Features analyzed include age, sex, cancer stage, grade, and various clinical factors. We excluded features with over 70% missing data, patients without follow-up times, duplicates, and those with a survival time of zero. The SEER cohort is available for download at www.seer.cancer.gov.
E.2 Survival learners
In this section, we describe the implementation and hyperparameters of the survival learners.
CoxPH: The Cox Proportional Hazards model (CoxPH) is a semi-parametric survival analysis model [9], consisting of a non-parametric baseline hazard function $h_{0}(t)$ and a parametric partial hazard function $f(\bm{\theta},\bm{x}_{i})$, typically estimated using the Breslow method [48]. The hazard function is $h(t\mid\bm{x}_{i})=h_{0}(t)\exp(f(\bm{\theta},\bm{x}_{i}))$, where $f$ is linear. The MLE of $\hat{\bm{\theta}}$ maximizes the Cox partial log-likelihood [9]. CoxPH is implemented from the scikit-survival package version 0.21.0.
GBSA: GBSA [39] is a likelihood-based boosting model that optimizes an $\ell_{2}$-norm penalized Cox partial log-likelihood [9]. It updates candidate variables flexibly at each boosting step using an offset-based gradient boosting method, unlike traditional gradient boosting, which updates either a single component or all features for fitting the gradient. GBSA is implemented from the scikit-survival package version 0.21.0.
RSF: Random Survival Forests (RSF) [15] extend decision trees for survival analysis by constructing multiple trees on bootstrapped samples. Each tree predicts an individual’s survival function, with splits based on criteria like the log-rank test. RSF handles censored data as right-censored and accounts for it during tree construction, with final predictions as the average of all trees’ survival functions. RSF is implemented from the scikit-survival package version 0.21.0.
DeepSurv: DeepSurv [13] is a neural network implementation of the CoxPH model of the form $h\left(t\mid\bm{x}_{i}\right)=h_{0}\left(t\right)\exp\left(f\left(\bm{\theta},\bm{x}_{i}\right)\right)$, where $f\left(\bm{\theta},\bm{x}_{i}\right)$ denotes a risk score as a nonlinear function of the features, i.e.$f\left(\bm{\theta},\bm{x}_{i}\right)=\sigma(\bm{x}_{i}\bm{\theta})$, where $\sigma$ is a nonlinear activation function. Popular functions include the hyperbolic tangent (Tanh) function or the Rectified Linear Unit (ReLU). The MLE for $\hat{\bm{\theta}}$ is derived by numerically maximizing the partial Cox log-likelihood. DeepSurv is implemented from https://github.com/shi-ang/BNN-ISD/.
MTLR: Multi-Task Logistic Regression (MTLR) [16] is a neural network implementation of a discrete-time model that predicts the survival distribution using a sequence of dependent logistic regressors. The number of discrete times is determined by the square root of the number of uncensored patients, and we use quantiles to divide those uncensored instances evenly into each time interval. MTLR is implemented from https://github.com/shi-ang/BNN-ISD/.
Table 5 reports the selected hyperparameters for the survival learners.
Table 5: Hyperparameters for the survival learners.
Model Parameter Value
CoxPH alpha 0.0001
ties ’Breslow’
n_iter 100
tol 1e-9
GBSA n_estimators 100
max_depth 1
min_samples_split 2
min_samples_leaf 1
max_features ’sqrt’
sub_sample 0.8
random_state 0
RSF n_estimators 100
max_depth 1
min_samples_split 2
min_samples_leaf 1
max_features ’sqrt’
random_state 0
DeepSurv hidden_size 100
verbose False
lr 0.001
c1 0.01
num_epochs 1000
batch_size 32
dropout 0.25
early_stop True
patience 10
MTLR verbose False
lr 0.001
c1 0.01
num_epochs 1000
batch_size 32
early_stop True
patience 10
E.3 Implementation details
Copula model. To validate our approach in semi-synthetic datasets, we train two copula models, Clayton and Frank, alongside two nonlinear Weibull proportional-hazards models for the marginal distributions $T$ and $C$, respectively. The Weibull models use a multilayer perceptron (MLP) as the backbone network architecture. We use a single layer with 32 hidden nodes and the ReLU activation function. For all experiments, we train the models for 10000 epochs using the Adam optimizer [49] with a learning rate of 1e-3, batch normalization enabled, and a batch size equal to the number of training instances. We use early stopping to terminate the training process if the validation loss does not improve for 100 consecutive epochs.
Loss function. We estimate the model parameters using maximum likelihood estimation (MLE) under dependent censoring, as described in [17]. Given survival data $\mathcal{D}$, let $f_{T\mid\bm{x}}$ and $F_{T\mid\bm{x}}$ represent the conditional density and cumulative distribution functions, respectively, over the event horizon, with the event distribution $T$ and the censoring distribution ${C}$. Under independence censoring, the likelihood of the $i$-th instance experiencing the event is:
$\mathcal{L}(\mathcal{D})=\prod_{i=1}^{N}\left[\text{Pr}(T=t_{i},C\geq t_{i}\mid\bm{x}=\bm{x}_{i})^{\delta_{i}}\times\text{Pr}(T\geq t_{i},{C}=t_{i}\mid\bm{x}=\bm{x}_{i})^{1-\delta_{i}}\right].$ (E.3)
However, when $T$ and $C$ are no longer independent, we can no longer rely on this clean decomposition of the survival likelihood. Based on Sklar’s theorem [42], in order to adjust Equation E.3 to dependent censoring under an Archimedean copula, we introduce the following lemma [17]:
Lemma 1.
Under Sklar’s (Survival) Theorem, and given $C(u_{1},u_{2})$ is a Archimedean copula, then if
$S_{T,C}(t,c|\bm{x})=C(u_{1},u_{2})\Big{|}^{u_{1}=S_{T}(t\rvert\bm{x})}_{u_{2}=S_{C}(c\rvert\bm{x})},$ (E.4)
then we have
$\int_{c}^{\infty}f_{C}(c\mid t,\bm{x})=\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\Big{|}^{u_{1}=S_{T}(t\rvert\bm{x})}_{u_{2}=S_{C}(c\rvert\bm{x})}.$ (E.5)
Applying Lemma 1 to Equation E.3 and taking the sum of logs yields the log-likelihood for survival models under dependent censoring:
$\mathcal{L}(\mathcal{D})=\sum_{i=1}^{N}\delta_{i}\;\text{log}\left[f_{T}(t_{i}|\bm{x}_{i})\right]\;+\;\delta_{i}\;\text{log}\left[\frac{\partial}{\partial u_{1}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T}(t_{i}\rvert\bm{x}_{i})}_{u_{2}=S_{C}(t_{i}\rvert\bm{x}_{i})}\right]$
$+\;(1-\delta_{i})\;\text{log}\left[f_{C}(t_{i}|\bm{x}_{i})\right]+(1-\delta_{i})\;\text{log}\left[\frac{\partial}{\partial u_{2}}C(u_{1},u_{2})\rvert^{u_{1}=S_{T\rvert}(t_{i}\rvert\bm{x}_{i})}_{u_{2}=S_{C}(t_{i}\rvert\bm{x}_{i})}\right].$ (E.6)
In this expression, the first term represents the log-likelihood of observing the event at time $t_{i}$. The second term captures the conditional probability of observing the censoring time after the event time, given that the event occurred at $t_{i}$. The third and fourth terms, by symmetry, represent the same quantities for the censoring time. Given $C(u_{1},u_{2})$ is an Archimedean copula (e.g., Clayton or Frank), the partial derivatives in Equation E.3 have closed-form solutions, allowing the log-likelihood to be expressed in a closed form that can be optimized using gradient-based methods. To estimate Equation E.3, we follow the approach described by Foomani et al. [17], as outlined in Algorithm 1. The parameters are optimized using gradient descent with the Adam optimizer [49].
Algorithm 1 Learning Under Dependent Censoring. Reproduced from [17].
1:Dataset $\mathcal{D}=\left\{\left(\bm{x}_{i},t_{i},\delta_{i}\right)\right\}_{i=1}^{N}$; Copula $C_{\theta}$: bivariate copula parameterized by $\theta$; $M$: survival model class parameterized by $\phi$ that can produce $\hat{S}^{(M)}_{T}(t|\bm{x})$ and $\hat{f}^{(M)}_{T}(t|\bm{x})$; $\alpha$: learning rate; $M$: number of training epochs; $K$: large constant; $\theta_{\min}$: small positive threshold for $\theta$.
2:$\hat{\theta},\hat{\phi}_{T},\hat{\phi}_{C}$: learned parameters of the copula and each marginal survival model.
3:$M_{T}\leftarrow\text{Instantiate}(M;\hat{\psi}^{(0)}_{T})$
4:$M_{C}\leftarrow\text{Instantiate}(M;\hat{\psi}^{(0)}_{C})$
5:$C_{\theta}\leftarrow\text{Instantiate}(C;\hat{\theta}^{(0)})$
6:for $i=1$ to $M$ do
7:     $L_{i}\leftarrow\ell(D;\hat{f}^{(M_{T})}_{T},\hat{f}^{(M_{C})}_{T},\hat{S}^{(M_{T})}_{T},\hat{S}^{(M_{C})}_{T},C_{\hat{\theta}^{(i)}})$
8:     $\hat{\psi}_{C}^{(i)}\leftarrow\text{Adam}(L_{i},\hat{\psi}_{C},\alpha)$
9:     $\hat{\psi}_{T}^{(i)}\leftarrow\text{Adam}(L_{i},\hat{\psi}_{T},\alpha)$
10:     $\nabla\hat{\theta}^{(i)}\leftarrow\nabla\hat{\theta}^{(i)}\times K$
11:     $\nabla\hat{\theta}^{(i)}\leftarrow\nabla\hat{\theta}^{(i)}\big{|}_{[-0.1,0.1]}$
12:     $\hat{\theta}^{(i)}\leftarrow\text{Adam}(L_{i},\hat{\theta},\alpha)$
13:     $\hat{\theta}^{(i)}\leftarrow\min(\hat{\theta}^{(i)},\theta_{\min})$
14:end for
15:return $\hat{\theta}^{(i)},\hat{\psi}_{T}^{(i)},\hat{\psi}_{C}^{(i)}$
Reproducibility. All experiments were implemented and executed in Python 3.9 with PyTorch 1.13.1, NumPy 1.24.3, and Pandas 1.5.3. The Copula-Graphic estimator was implemented using the compound.Cox package version 3.32. Tensor computations were carried out in double precision (fp64). The experiments were conducted on a single workstation equipped with an Intel Core i9-10980XE 3.00GHz CPU, 64GB of RAM, and an NVIDIA GeForce RTX 3090 GPU. Model training was performed on the GPU using CUDA 11.7. All datasets utilized in this study are publicly available. Detailed instructions for reproducing the results are included in the supplementary code.
Appendix F Complete results
F.1 Synthetic data
Dependent Metrics Under Misspecification. Since our metrics require the practitioners to specify the Archimedean copula family and dependency level a priori, we find it important to evaluate our metrics under misspecification – i.e., where either the copula family or the dependency level is incorrectly specified. To this end, we implement three forms of misspecification: (1) The copula family is misspecified (i.e., we assume the family is Clayton [10], when in reality, the data comes from a Frank [24] copula, and vice-versa), but the dependency strength (Kendall’s $\tau$) is correct. (2) The dependency strength is misspecified (i.e., we invert the dependency strength relative to a fixed baseline of Kendall’s $\tau$ = 0.8, so the “wrong” Kendall’s $\tau$ becomes 0.8 - Kendall’s $\tau$). (3) Both the copula family and the dependency strength are misspecified given the aforementioned procedure.
Figure 5 shows the evaluation bias when only the copula family is misspecified (Option 1). We see that this form of misspecification leads to significantly increased error margins for the dependent metrics, however, they are still very much on par with their independent variants. Figure 6 shows the evaluation bias when only the dependency strength is misspecified (Option 2). As before, the proposed metrics exhibit increased error margins, but the gap between independent and dependent metrics narrows further. In fact, their performance is nearly identical in this setting, suggesting that misspecifying the dependence strength may have a limited impact on evaluation outcomes. Figure 7 shows the evaluation bias when both the copula family and dependency strength are misspecified (Option 3). The differences between the metrics are again minor, but now the error margins of the dependent metrics almost overlap with those of the independent metrics. Under this kind of misspecification, the two approaches give nearly identical outcomes.
Dependent Metrics With Increasing Sample Size. We want to investigate whether sample size impacts the performance of metrics. Using a fixed Kendall’s $\tau$ = 0.5, two copula families (Clayton and Frank), and 10 features, we generate synthetic survival datasets with sample sizes ranging from 500 to 10,000. We again compare each metric’s estimate to its ground-truth value (computed using uncensored event times) and record the absolute error. Figure 8 shows the result. Regarding discrimination, Harrell’s CI occasionally gives better estimates than CI-Dep in smaller sample sizes, however, CI-Dep prevails in larger datasets, but with overlapping error margins. CI-Dep generally seems to outperform Uno’s CI, most but dominantly in larger sample sizes. The IBS-Dep and MAE-Dep give slightly better performances than their independent counterparts in large datasets, however, the differences between them are minor in low sample settings.
Figure 5: Plot of mean biases ($\pm$ 2 SD.) as a function of the dependency, where only the copula family is misspecified, averaged over 100 experiments ($N=10,000$, $d=10$).
Figure 6: Plot of mean biases ($\pm$ 2 SD.) as a function of the dependency, where only the dependency strength is misspecified, averaged over 100 experiments ($N=10,000$, $d=10$).
Figure 7: Plot of mean biases ($\pm$ 2 SD.) as a function of the dependency, where the copula family and the dependency strength are misspecified, averaged over 100 experiments ($N=10,000$, $d=10$).
Figure 8: Plot of mean biases ($\pm$ 2 SD.) as a function of the sample size generated from the DGP (from 500 to 10,000), averaged over 100 experiments (Kendall’s $\tau$ = 0.5, $d=10$).
F.2 Semi-synthetic data
Approximated dependency. Table 6 shows the dependence structure estimates for the semi-synthetic datasets in terms of Kendall’s $\tau$, optimized using MLE (see Equation E.3). The Kendall’s $\tau$ values indicate the strength and direction of dependence between variables, reflecting the association between event or survival times within each dataset. The results show that datasets like METABRIC and MIMIC-IV (all) exhibit stronger dependence, with Kendall’s $\tau$ reaching up to 0.36, whereas SEER datasets tend to have weaker associations, often close to zero. However, the censoring rates in these datasets is still generally low.
Discrimination. Table 7 shows the discrimination (CI) results in the 6 semi-synthetic datasets. The results show that the CI-Dep metric generally performs comparably to Uno’s C-index across datasets, with both showing slightly higher bias than Harrell’s C-index in most scenarios. While CI-Dep maintains stable performance in low-censoring settings like METABRIC, it exhibits increased sensitivity in high-censoring environments, as seen in the MIMIC-IV hospital dataset. For the larger SEER datasets, CI-Dep remains competitive with the baseline metrics, though feature selection strategies – particularly Top 5 – can degrade the performance.
Accuracy. Table 8 shows the accuracy (IBS and MAE) results in the 6 semi-synthetic datasets. Across all datasets, the proposed IBS-Dep and MAE-Dep metrics produced results similar to or slightly better than their independent counterparts. For instance, in METABRIC, the IBS-Dep remained consistent across all feature strategies, and the proposed MAE-Dep exhibited marginal improvements over the pseudo-observation and margin-based approaches. In MIMIC-IV, while the IBS values slightly increased under Top 5 and Top 10 strategies, the MAE-Dep maintained comparable bias to baseline metrics. Notably, in high-censoring scenarios such as MIMIC-IV (hospital), the proposed dependent metrics yielded more stable and interpretable estimates compared to IPCW, which showed higher variability. In SEER datasets (brain, liver, and stomach), the proposed metrics remained consistent across different feature strategies, with MAE-Dep often outperforming Hinge and Pseudo metrics in terms of lower bias.
Table 6: Dependence structure estimates in terms of Kendall’s $\tau$ by dataset and strategy.
Dataset Strategy Copula $\hat{\tau}$
METABRIC ($N$=1,102, $C$=3.8%) Original Frank 0.20
Top 5 Frank 0.18
Top 10 Frank 0.20
Rand. 25% Frank 0.07
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original Frank 0.22
Top 5 Clayton 0.31
Top 10 Clayton 0.36
Rand. 25% Frank 0.16
MIMIC-IV (hospital) ($N$=6,780, $C$=74.3%) Original Frank 0.06
Top 5 Frank 0.04
Top 10 Frank 0.05
Rand. 25% Frank 0.07
SEER (brain) ($N$=44,137, $C$=9.0%) Original Frank 0.03
Top 5 Frank 0.04
Top 10 Frank 0.03
Rand. 25% Frank 0.08
SEER (liver) ($N$=51,704, $C$=20.5%) Original Frank 0.04
Top 5 Frank 0.01
Top 10 Clayton 0.01
Rand. 25% Frank 0.00
SEER (stomach) ($N$=56,807, $C$=21.9%) Original Frank 0.00
Top 5 Clayton 0.05
Top 10 Frank 0.01
Rand. 25% Frank 0.00
Table 7: Comparison of baseline and proposed discrimination metrics in terms of the mean ($\pm$ 2 SD.) bias, i.e., the proximity to the true error, averaged over 5 survival learners and 5 experiments.
Dataset Strategy Harrell Uno Dep.
METABRIC ($N$=1,102, $C$=3.8%) Original 0.001$\pm$0.002 0.004$\pm$0.003 0.005$\pm$0.005
Top 5 0.001$\pm$0.002 0.005$\pm$0.003 0.006$\pm$0.005
Top 10 0.001$\pm$0.002 0.004$\pm$0.002 0.005$\pm$0.003
Rand. 25% 0.001$\pm$0.002 0.023$\pm$0.014 0.023$\pm$0.014
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original 0.006$\pm$0.004 0.007$\pm$0.004 0.007$\pm$0.004
Top 5 0.011$\pm$0.004 0.016$\pm$0.005 0.016$\pm$0.005
Top 10 0.010$\pm$0.003 0.013$\pm$0.004 0.013$\pm$0.004
Rand. 25% 0.006$\pm$0.002 0.006$\pm$0.003 0.006$\pm$0.003
MIMIC-IV (hospital) ($N$=6,780, $C$=74.3%) Original 0.007$\pm$0.009 0.013$\pm$0.017 0.025$\pm$0.032
Top 5 0.008$\pm$0.011 0.012$\pm$0.011 0.038$\pm$0.043
Top 10 0.007$\pm$0.009 0.011$\pm$0.013 0.035$\pm$0.040
Rand. 25% 0.014$\pm$0.012 0.018$\pm$0.020 0.027$\pm$0.029
SEER (brain) ($N$=44,137, $C$=9.0%) Original 0.005$\pm$0.001 0.009$\pm$0.002 0.005$\pm$0.003
Top 5 0.003$\pm$0.001 0.009$\pm$0.003 0.006$\pm$0.003
Top 10 0.005$\pm$0.001 0.009$\pm$0.002 0.006$\pm$0.002
Rand. 25% 0.018$\pm$0.046 0.020$\pm$0.018 0.019$\pm$0.018
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.008$\pm$0.002 0.006$\pm$0.003 0.008$\pm$0.003
Top 5 0.019$\pm$0.002 0.017$\pm$0.003 0.017$\pm$0.003
Top 10 0.008$\pm$0.002 0.009$\pm$0.002 0.011$\pm$0.001
Rand. 25% 0.011$\pm$0.014 0.019$\pm$0.017 0.018$\pm$0.017
SEER (stomach) ($N$=56,807, $C$=21.9%) Original 0.004$\pm$0.002 0.003$\pm$0.002 0.005$\pm$0.002
Top 5 0.009$\pm$0.001 0.006$\pm$0.002 0.007$\pm$0.001
Top 10 0.004$\pm$0.002 0.003$\pm$0.002 0.004$\pm$0.002
Rand. 25% 0.006$\pm$0.006 0.014$\pm$0.012 0.013$\pm$0.012
Table 8: Comparison of baseline and proposed accuracy metrics (IBS and MAE) in terms of mean ($\pm$ 2 SD.) bias across 5 survival learners and 5 experiments. KM denotes the proposed IBS metric under the independent copula for comparison.
IBS MAE
Dataset Strategy IPCW KM Dep. Hinge Pseudo Margin Dep.
METABRIC ($N$=1,102, $C$=3.8%) Original 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.246$\pm$0.414 0.314$\pm$0.459 0.328$\pm$0.431
Top 5 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.249$\pm$0.398 0.308$\pm$0.460 0.298$\pm$0.464
Top 10 0.012$\pm$0.017 0.012$\pm$0.017 0.012$\pm$0.017 1.397$\pm$0.416 0.248$\pm$0.418 0.317$\pm$0.462 0.308$\pm$0.465
Rand. 25% 0.012$\pm$0.018 0.012$\pm$0.018 0.012$\pm$0.018 1.397$\pm$0.416 0.266$\pm$0.435 0.342$\pm$0.463 0.328$\pm$0.467
MIMIC-IV (all) ($N$=12,845, $C$=10.6%) Original 0.008$\pm$0.003 0.009$\pm$0.003 0.009$\pm$0.003 62.244$\pm$10.539 37.156$\pm$10.255 37.156$\pm$10.255 37.301$\pm$10.245
Top 5 0.008$\pm$0.004 0.011$\pm$0.004 0.011$\pm$0.004 65.626$\pm$11.943 39.558$\pm$11.741 39.558$\pm$11.741 39.717$\pm$11.743
Top 10 0.008$\pm$0.004 0.010$\pm$0.004 0.010$\pm$0.004 63.716$\pm$11.380 38.131$\pm$11.190 38.131$\pm$11.190 38.312$\pm$11.205
Rand. 25% 0.008$\pm$0.003 0.011$\pm$0.004 0.011$\pm$0.004 62.749$\pm$8.117 37.204$\pm$8.102 37.204$\pm$8.102 37.364$\pm$8.062
MIMIC-IV (hospital) ($N$=6,780, $C$=74.3%) Original 0.088$\pm$0.030 0.066$\pm$0.029 0.066$\pm$0.029 6.438$\pm$0.548 3.107$\pm$0.548 4.542$\pm$0.528 4.302$\pm$0.490
Top 5 0.088$\pm$0.030 0.064$\pm$0.029 0.064$\pm$0.029 6.447$\pm$0.527 3.098$\pm$0.521 4.646$\pm$0.501 4.293$\pm$0.808
Top 10 0.087$\pm$0.030 0.066$\pm$0.029 0.066$\pm$0.029 6.446$\pm$0.540 3.112$\pm$0.545 4.603$\pm$0.569 4.250$\pm$0.857
Rand. 25% 0.104$\pm$0.015 0.079$\pm$0.018 0.079$\pm$0.018 6.513$\pm$0.383 3.057$\pm$0.379 4.791$\pm$0.386 4.617$\pm$0.391
SEER (brain) ($N$=44,137, $C$=9.0%) Original 0.004$\pm$0.003 0.004$\pm$0.007 0.004$\pm$0.007 1.208$\pm$0.076 1.501$\pm$0.122 1.113$\pm$0.085 1.132$\pm$0.086
Top 5 0.004$\pm$0.003 0.004$\pm$0.007 0.004$\pm$0.007 1.224$\pm$0.077 1.495$\pm$0.118 1.106$\pm$0.084 1.126$\pm$0.084
Top 10 0.004$\pm$0.003 0.004$\pm$0.007 0.004$\pm$0.007 1.212$\pm$0.076 1.497$\pm$0.120 1.108$\pm$0.084 1.128$\pm$0.086
Rand. 25% 0.005$\pm$0.002 0.002$\pm$0.002 0.002$\pm$0.002 1.288$\pm$0.119 1.463$\pm$0.098 1.064$\pm$0.114 1.080$\pm$0.107
SEER (liver) ($N$=51,704, $C$=20.5%) Original 0.003$\pm$0.003 0.003$\pm$0.002 0.003$\pm$0.002 2.498$\pm$0.075 2.475$\pm$0.134 1.478$\pm$0.110 1.520$\pm$0.107
Top 5 0.003$\pm$0.003 0.003$\pm$0.002 0.003$\pm$0.002 2.688$\pm$0.122 2.246$\pm$0.134 1.250$\pm$0.123 1.226$\pm$0.219
Top 10 0.003$\pm$0.003 0.002$\pm$0.002 0.002$\pm$0.002 2.513$\pm$0.072 2.462$\pm$0.119 1.466$\pm$0.102 1.450$\pm$0.170
Rand. 25% 0.005$\pm$0.005 0.004$\pm$0.004 0.004$\pm$0.004 2.505$\pm$0.136 2.499$\pm$0.272 1.490$\pm$0.193 1.524$\pm$0.227
SEER (stomach) ($N$=56,807, $C$=21.9%) Original 0.003$\pm$0.002 0.002$\pm$0.002 0.002$\pm$0.002 2.119$\pm$0.066 2.189$\pm$0.189 1.246$\pm$0.119 1.278$\pm$0.117
Top 5 0.003$\pm$0.002 0.002$\pm$0.002 0.002$\pm$0.002 2.145$\pm$0.071 2.166$\pm$0.193 1.223$\pm$0.124 1.185$\pm$0.210
Top 10 0.003$\pm$0.002 0.002$\pm$0.002 0.002$\pm$0.002 2.106$\pm$0.070 2.207$\pm$0.193 1.264$\pm$0.126 1.250$\pm$0.221
Rand. 25% 0.003$\pm$0.001 0.002$\pm$0.001 0.002$\pm$0.001 2.090$\pm$0.115 2.239$\pm$0.176 1.325$\pm$0.105 1.311$\pm$0.202
Ranking. In general, a (censored) metric should be representative of the true metric, i.e., it should tell us which model makes the best predictions. Therefore, we create a new experiment as follows:
  • 1.
    In our semi-synthetic datasets, we are aware of the ground-truth (the actual event times) of all instances. For each model, we use this information to calculate the True CI, True IBS, and True MAE results – that is, how good are the survival predictions using each metric, if all the instances were observed (not censored).
  • 2.
    For each true metric, we then rank the top-3 survival models – e.g., CoxPH, RSF, and DeepSurv – by their true metric.
  • 3.
    For each censored metric (proposed and existing), we check if the metric gives the same ordering of survival models as its corresponding true metric.
Figures 9-14 show the ranking accuracy, i.e., the number of times the resp. metric correctly (green dot) or incorrectly (red dot) identified the ordering of the top-3 survival learners by the true metric, over 5 experiments in 6 datasets. We see general alignment between the metrics, but IBS-Dep give notable improvements over its independent variant in the SEER (liver) and SEER (stomach) datasets.
Figure 9: Ranking accuracy over 5 experiments in the METABRIC dataset.
Figure 10: Ranking accuracy over 5 experiments in the MIMIC-IV (all) dataset.
Figure 11: Ranking accuracy over 5 experiments in the MIMIC-IV (hospital) dataset.
Figure 12: Ranking accuracy over 5 experiments in the SEER (brain) dataset.
Figure 13: Ranking accuracy over 5 experiments in the SEER (liver) dataset.
Figure 14: Ranking accuracy over 5 experiments in the SEER (stomach) dataset.
Appendix G Derivations
G.1 The Kaplan-Meier estimator
This derivation is reproduced from [4, Ch. 2].
Recall that $(t_{i},\delta_{i}),\;i=1,\ldots,n$ represents survival data without covariates, where $t_{i}=\min(e_{i},c_{i})$, $\delta_{i}=\mathbbm{1}[e_{i}\leq c_{i}]$. We can then consider the survival function $S(t)$ as a function that decreases in steps, with discontinuities occurring only at the observed times of death. In this case, the survival function for the event can be expressed as
$S_{T}(t)=\text{Pr}(T\geq t)=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{\text{Pr}(T=t_{i})}{\text{Pr}(T\geq t_{i})}\right).$ (G.1)
Second, suppose that $T$ and $C$ are independent. Then, we can write
$S_{T}(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(T=t_{i},C\geq t_{i})\right)\text{Pr}(T\geq t_{i},C\geq t_{i})$ (G.2)
$=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\text{Pr}(\min(T,C)=t_{i},T\leq C)\right)\text{Pr}(\min(T,C)\geq t_{i}).$ (G.3)
Finally, we replace the probability ratio of the last expression by its estimate to obtain
$S_{T}(t)$ $=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{\sum_{\ell=1}^{N}I(t_{\ell}=t_{i},\delta_{\ell}=1)/n}{\sum_{\ell=1}^{N}I(t_{\ell}\geq t_{i})/n}\right)$ (G.4)
$=\prod_{t_{i}\leq t,\,d_{i}=1}\left(1-\frac{1}{n_{i}}\right).$ (G.5)
G.2 The Copula-Graphic estimator
This derivation is reproduced from [4, Ch. 4].
Consider an Archimedean copula model
$P(T>t,C>t)=\phi_{\theta}^{-1}\big{[}\phi_{\theta}\{S_{T}(t)\}+\phi_{\theta}\{S_{C}(t)\}\big{]},$ (G.6)
where $\phi_{\theta}:[0,1]^{d}\rightarrow[0,\infty]$ is a generator function, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [4, Ch. 3], and $S_{T}(t)=\text{Pr}(T>t)$ and $S_{C}(t)=\text{Pr}(C>t)$ are the marginal survival functions. Assume that $S_{T}(t)$ is a decreasing step function with jumps at event times. Thus, $\delta_{i}=1$ implies $S_{T}(t_{i})\neq S_{T}(t_{i}-dt)$ and $S_{C}(t_{i})=S_{C}(t_{i}-dt)$. Setting $t=t_{i}$ in Equation G.6, we have
$\text{Pr}(T\geq t_{i},C\geq t_{i})=\phi_{\theta}\{S_{T}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i})\}.$ (G.7)
In the left-side of the preceding equation, we estimate $\text{Pr}(T>t_{i},C>t_{i})$ by ($n_{i}-1)/n$, where $n_{i}-1=\sum_{\ell=1}^{n}\mathbbm{1}[t_{\ell}>t_{i}]$ is the number of event-free instances at time $t_{i}$. Then we have
$\phi_{\theta}\left(\frac{n_{i}-1}{n}\right)=\phi_{\theta}\{S_{T}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i})\}.$ (G.8)
We set $t=c=t_{i}-dt$ in Equation G.6 and then estimate $\text{Pr}(T>t_{i}-dt,C>t_{i}-dt)$ by $n_{i}/n$. Then,
$\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{T}(t_{i}-dt)\}+\phi_{\theta}\{S_{C}(t_{i})\},\quad\delta_{i}=1$ (G.9)
Equations G.8 and G.9 result in the system of difference equations
$\phi_{\theta}(\frac{n_{i}-1}{n})-\phi_{\theta}(\frac{n_{i}}{n})=\phi_{\theta}\{S_{T}(t_{i})\}+\phi_{\theta}\{S_{C}(t_{i}-dt)\},\quad\delta_{i}=1$ (G.10)
We apply the standard constraint that $S_{T}(t_{i}-dt)=1$ when $t_{i}$ is the earliest event time. Under this condition, the solution to the differential equations is
$\phi_{\theta}\{S_{T}(t)\}$ $=\sum_{t_{i}\leq t,\delta_{i}=1}[\phi_{\theta}\{S_{T}(t_{i})\}-\phi_{\delta}\{S_{T}(t_{i}-dt)\}]$ (G.11)
$=\sum_{t_{i}\leq t,\delta_{i}=1}\phi_{\delta}\left(\frac{n_{i}-1}{n}\right)-\phi_{\delta}\left(\frac{n_{i}}{n}\right),$ (G.12)
which is equivalent to the CG estimator. Under the independence copula, given by $\phi_{\delta}(t)=-\log(t)$, the CG estimator is equivalent to the Kaplan-Meier estimator. Substituting for the Clayton copula [10]
$S_{T}(t)=\left[1+\sum_{t_{i}\leq t,\delta_{i}=1}\left\{\left(\frac{n_{i}-1}{n}\right)^{-\theta}-\left(\frac{n_{i}}{n}\right)^{-\theta}\right\}\right]^{-1/\theta}.$ (G.13)